### v2

# EnScale: Temporally-consistent multivariate generative downscaling via proper scoring rules

Abstract
The practical use of future climate projections from global circulation models (GCMs) is often limited by their coarse spatial resolution, requiring downscaling to generate high-resolution data. Regional climate models (RCMs) provide this refinement, but are computationally expensive. To address this issue, machine learning models can learn the downscaling function, mapping coarse GCM outputs to high-resolution fields. Among these, generative approaches aim to capture the full conditional distribution of RCM data given coarse-scale GCM data, which is characterized by large variability and thus challenging to model accurately. We introduce EnScale, a generative machine learning framework emulating the full GCM-to-RCM map by training on multiple pairs of GCM and corresponding RCM data. It first adjusts large-scale mismatches between GCM and coarsened RCM data, followed by a super-resolution step to generate high-resolution fields. To efficiently model the high-dimensional output, the super-resolution step employs a novel class of sparse local stochastic layers. Both steps employ generative models optimized with the energy score, a proper scoring rule. Compared to state-of-the-art ML downscaling approaches, our setup reduces computational cost by about one order of magnitude. EnScale jointly emulates multiple variables – temperature, precipitation, solar radiation, and wind – spatially consistent over Central Europe. In addition, we propose a variant EnScale-t that enables temporally consistent downscaling. We establish a comprehensive evaluation framework across various categories including calibration, spatial structure, extremes, and multivariate dependencies. Comparison with diverse benchmarks demonstrates EnScale’s strong performance and computational efficiency. EnScale offers a promising approach for accurate and temporally consistent RCM emulation.
1 Introduction
General circulation models (GCMs) provide us with future climate projections that enable assessing potential changes in climate under different greenhouse gas emission scenarios. Multiple research groups around the world develop and maintain GCMs that incorporate different modeling assumptions and parameterizations to simulate the climate system of the Earth. By virtue of international efforts such as the Coupled Model Intercomparison Project (CMIP) [Eyring2016OverviewOrganization, Taylor2012AnDesign], large ensembles of GCM data are available for various Representative Concentration Pathway (RCP) / Shared Socioeconomic Pathway (SSP) scenarios.
However, these physics-based models are typically computed at a coarse spatial resolution. GCMs are able to simulate large-scale climate patterns, but often lack the spatial resolution needed to capture climate variations at impact-relevant regional and local scales [Benestad2010DownscalingExtremes, Maraun2016BiasReview, Rampal2024EnhancingLearning, Wilby1997DownscalingLimitations]. Downscaling describes the process of increasing the spatial and/or temporal resolution of GCM outputs. The common approach of dynamical downscaling relies on process-based regional climate models (RCMs) that use GCM data as boundary conditions, i.e., RCMs are “driven” by the corresponding GCM. They refine the GCM’s large-scale climate information and yield high-resolution projections [Giorgi2019ThirtyNext]. Due to necessary simplifications and parameterizations, regional climate models differ in their representation of processes, resulting in different RCMs, and initiatives such as the Coordinated Regional Climate Downscaling Experiment (CORDEX) [Giorgi2015RegionalInitiative] provide data from multiple models for several regions around the globe.
However, RCMs remain computationally costly. To address this challenge, statistical downscaling establishes a statistical relationship between low-resolution (LR) and high-resolution (HR) data. In addition to more traditional approaches (e.g. [Maraun2010PrecipitationUser, Maraun2018StatisticalResearch]), machine learning (ML) for resolution enhancement has been applied to predict high-resolution outputs from coarse data in various contexts. This includes downscaling GCMs to RCM resolutions [Rampal2024EnhancingLearning], targeting convection-permitting models [Addison2022] or high-resolution observational data [Harris2022AForecasts, Leinonen2020, PriceStephanRasp2022IncreasingModels, Vosper2023DeepScales], often using varying sets of predictors.
Many different setups for ML-based downscaling have been considered. Among these, some attempts directly approximate the full path from GCM to RCM data: the LR predictors stem from GCM data and the predictands from the RCM [Bano-Medina2021]. We refer to this as RCM emulation. Since there are modeling biases between global and regional outputs, downscaling GCMs to RCMs requires learning a transformation map in addition to the pure resolution refinement [vanderMeer2023DeepFrameworks]. Similar mismatches can arise in other setups, for example with predictors from a biased numerical weather model and targets from radar measurements [Harris2022AForecasts, PriceStephanRasp2022IncreasingModels]. Alternative setups include super-resolution and perfect prognosis downscaling. In the super-resolution setting, the inputs for the ML model are artificially coarsened versions of the HR targets, so that the predictor is exactly a degraded form of the target field [Harder2023Hard-ConstrainedDownscaling, Leinonen2020, Stengel2020, Vosper2023DeepScales]. Other works instead use different climate variables as predictors, e.g. by learning the relationship between large-scale atmospheric predictors and a local HR target. In perfect prognosis downscaling, this relationship is learned from observational data [Bano-Medina2020ConfigurationDownscaling], whereas [Doury2022] and [Doury2024OnPrecipitation] learn it from RCM data. Note that they also refer to this task as “RCM emulation”, which is different from our RCM emulation: they consider both predictors and targets from the RCM, whereas we use predictors from the GCM and targets from the RCM. Some studies conclude that emulators can successfully learn the GCM-RCM map directly [vanderMeer2023DeepFrameworks], whereas others find better performance for the pure super-resolution task compared to a setup that additionally needs to learn large-scale shifts [Harris2022AForecasts]. In addition, learning the transformation map and the downscaling map separately can lead to improved results [PriceStephanRasp2022IncreasingModels].
Many downscaling approaches focus on producing a deterministic map at the fine resolution, given LR predictors [Bano-Medina2020ConfigurationDownscaling, Bano-Medina2021, Doury2022]. However, to capture the full variability, including extreme events, a stochastic approach is needed. In this case, the goal is to model the conditional distribution of the HR target given the LR input. Generative ML methods provide a way to achieve this by “learning” a complex, high-dimensional distribution from training data. After training, generative models can generate samples that reflect the variability present in the training distribution. Examples of generative models are Generative Adversarial Networks (GANs) [Goodfellow2020GenerativeNetworks] and diffusion models [Ho2020DenoisingModels, KarrasElucidatingModels, Song2019GenerativeDistribution, Song2020Score-BasedEquations]. In past studies, both approaches have been used successfull, e.g. to downscale precipitation [Addison2022, Harris2022AForecasts, Leinonen2020, Vosper2023DeepScales] or wind [Miralles2022, Stengel2020, White2019]. While GANs can often generate realistic samples, their optimization can be unstable [Arjovsky2017TowardsNetworks], and they are prone to mode collapse, where the model fails to represent the full diversity of the target distribution. Diffusion models, in contrast, are effective in capturing complex distributions, but are computationally costly, especially during inference.
Besides GANs and diffusion models, it is possible to train a generative model by minimizing a so-called proper scoring rule [Shen2024Engression:Regression, Alet2025SkillfulMarginals, Chen2022GenerativePost-processing, Kochkov2024NeuralClimate, Lang2024AIFS-CRPS:Score, Pacchiardi2024ProbabilisticMinimization]. Such a scoring rule compares the generated distribution by the machine learning model with the target RCM distribution. It is minimal if and only if the generated and target distribution agree. This loss function allows for easy optimization and fast training. In addition, it is flexible and can be combined with different model architectures.
Temporal consistency of downscaled data is important for accurate representation of multi-day statistics, for example in the assessment of heatwaves or accumulated precipitation. Yet, most existing methods do not consider temporal consistency. A few studies perform refinement in a temporal dimension, generating multiple time steps on a high-resolution scale for one or more input fields [Glawion2023SpateGAN:Approach, Serifi2021, Bassetti2024DiffESM:Models]. However, they still process each input patch separately of previous ones and, thus, do not ensure temporal consistency of longer time series on the high-resolution scale. In contrast, AI-based weather forecasting commonly employs an autoregressive approach: previous time steps are included as predictors, and an iterative application of the ML model can provide temporally consistent predictions.
In this study, we emulate the conditional distribution of the RCM data giving the corresponding GCM input. We train our model on a joint dataset combining multiple pairs of RCMs and GCMs. To make best use of this pooled dataset, we introduce a novel approach that splits the problem into several physically meaningful steps, similar to [PriceStephanRasp2022IncreasingModels]. We first correct for large-scale mismatches between the driving GCM and a coarsened version of the target RCM. Second, we perform super-resolution from the coarsened RCM to the high-resolution RCM. For this, we progressively refine the resolution in multiple stages [Shen2025ReverseDistributions]. In addition, to handle the high-dimensional output in a computationally efficient manner, we introduce a new architecture with sparse local stochastic layers. The generated values at a specific location only depend on the values at the lower-resolution in a local neighborhood. This architecture enables learning location-specific features while at the same time limiting the number of model parameters. For each step, we train a generative model minimizing the energy score, a multivariate generalization of the Continuous Ranked Probability Score (CRPS), one example of a proper scoring rule. Accordingly, we call our method EnScale (energy score + downscaling).
Our contributions are as follows:
  1. 1.
    We demonstrate that emulation of RCMs from GCM data in a multivariate and generative setup is feasible, both capturing the variability of RCM targets as well as multivariate dependencies by jointly emulating temperature (tas), precipitation (pr), solar radiation (rsds), and surface wind (sfcWind).
  2. 2.
    We propose new methods for RCM emulation: EnScale and EnScale-t, which combine the idea of downscaling via coarse correction with a novel generative model architecture and loss. They achieve great performance, while at the same time being computationally lightweight.
  3. 3.
    EnScale is our version that downscales each day independently of previous days, whereas EnScale-t conditions on the previous day and allows for generating temporally consistent time series.
  4. 4.
    We establish a comprehensive evaluation framework with metrics for overall performance, spatial and temporal structure, behavior in the extremes, and variable dependencies. For comparison, we propose a set of benchmarks, including physically motivated approaches and state-of-the-art generative models.
2 Setup and methods
2.1 Downscaling as a generative modeling problem
We view downscaling as a generative modeling problem as follows. Let $X$ denote the daily GCM data, consisting of multiple climate variables across all grid cells in Europe, with distribution $p_{X}$. Likewise, we call $Y$ the corresponding downscaled RCM data on our target area in Central Europe, with distribution $p_{Y}$. We denote all probability distributions by their corresponding density $p$. We aim to model the conditional distribution $p_{Y|X}$. Intuitively, this conditional $p_{Y|X}$ represents the distribution that we would obtain if we could re-run the same RCM with slightly perturbed GCM boundary conditions. The RCM output has much variability that is not present in the GCM input, as we discuss in Sec. 6. Physically, this can be interpreted as internal variability: different plausible realizations of local weather that are consistent with the same boundary forcing. Thus, multiple draws from the conditional distribution $p_{Y|X}$ correspond to such alternative realizations of internal variability. When these draws are produced by our ML model (or a baseline), we will refer to them as “samples”, or alternatively as “predictions”. Note that, in our case, “predictions” refer to the output of our downscaling models, not to forecasts of future climate (downscaled fields always correspond to the same time point as the GCM input). The predictions from our model can capture differences between RCMs, and we investigate various sources of variability later in Sec. 6.
Figure 1: Illustration of the dataset. The first row shows the GCM data from CNRM-CM5 on two example days with similar precipitation fields, displayed on the full spatial extent that is used as the model input. The second row presents the same GCM data cropped to approximately match the target area used in this study. The fourth row shows the corresponding target RCM data (ALADIN63 and RegCM4-6) that downscaled this GCM. The third row presents these RCM fields manually coarsened to a resolution comparable to the GCM data. Small differences between shape and domain of cropped GCM (row two) and coarsened RCM (row three) are due to small resolution differences and misaligned grids.
2.2 Downscaling with coarse correction
Our model is trained jointly on the entire dataset comprising multiple pairs of GCM data $X$ and corresponding RCM data $Y$. They stem from three different GCMs, and each driving GCM was downscaled by several RCMs (between two and three RCMs for one GCM, see Sec. 3). Despite training a single unified model, our goal is to capture both GCM- and RCM-specific characteristics. We aim to learn the conditional $p^{\iota}_{Y|X}$. Here, $\iota$ is an index for the GCM-RCM pair, emphasizing that we target the distribution for each GCM-RCM pair individually. Fig. 1 shows example GCM data on two selected days as well as corresponding downscaled data from two RCMs. These examples help to illustrate several challenges present in our heterogeneous dataset, as explained in the following.
Firstly, the GCM data differs from the RCM not only on local scales. Even when the RCM data is coarsened to a similar resolution as the GCM data, vast disagreements remain for some variables and days. For example, comparing the cropped GCM data with the pooled RCM data in Fig. 1 reveals distinct mismatches. Fig. 1 also illustrates the large variability of the conditional distribution of the coarsened RCM data given the GCM data. Even though the GCM input $X$ is similar for both days, the coarsened RCM data show considerable differences (see Sec. 6). These differences are due to internal variability of the climate system, as the GCM’s boundary conditions do not uniquely determine the weather conditions simulated by the RCM.
Secondly, each GCM-RCM pair follows a different distribution, and the ML model needs to learn the correct downscaling function for each of them. Fig. 1 compares the downscaled fields by two different RCMs, given the same driving GCM. We find that responses from multiple RCMs to the same GCM input differ considerably. Partly, this can be due to internal variability, as discussed above. In addition, there are disagreements between RCMs. This is because each RCM has parameterizations describing the effect of small-scale processes (e.g., clouds) that differ from those of other RCMs and the driving GCM. The land surface may be fixed in one model and respond to the atmosphere in another. Our experiments suggest that most differences between the RCMs are already present in the first step from GCM to coarsened RCM data, see App. A.1.
To solve these challenges, we split the problem into two parts, separating the correction on coarse scales and the super-resolution task [PriceStephanRasp2022IncreasingModels, Wan2024StatisticalModels], as visualized in Fig. 2. In the first step, we train a generative model to learn the conditional distribution of coarsened RCM data $Z$ given the GCM inputs $X$, $p^{\iota}_{Z|X}$. For this, the RCM data is artificially coarsened through average pooling over patches of size $16\times 16$, producing additional data $Z$ for each day. As the resolutions of $X$ ($\approx$250\text{\,}\mathrm{k}\mathrm{m}$\$) and $Z$ ($\approx$170\text{\,}\mathrm{k}\mathrm{m}$\$) are similar, predicting $Z$ from $X$ does not include any “downscaling”. Instead, this step captures large-scale mismatches between GCM and RCM. The output of this step is relatively low-dimensional (dimension $8\times 8$ per variable) and we call this the coarse model.
For the second step, we learn the conditional distribution $p^{\iota}_{Y|Z}$. This is a pure super-resolution problem, as $Z$ is the average-pooled version of $Y$, excluding the large-scale mismatches between $X$ and $Y$. Hence, we refer to the corresponding generative model as the super-resolution model. The transformation from $Z$ to $Y$ is high-dimensional, but the training data is more homogeneous as the different GCM-RCM pairs behave very similarly in this step (see App. A.1). We split the super-resolution model again into several resolution steps, each one learning a simpler conditional distribution (see Sec. 2.4).
In each step of our modeling chain, we add predictor variables indicating the GCM and RCM considered (see App. B.4 for how these predictors are integrated in the model architectures). This allows our model to learn the conditional $p^{\iota}_{Y|X}$ for each pair individually. It enables capturing model-specific details, while at the same time leveraging shared properties from all pairs, benefiting from a large dataset.
All steps are trained independently. We tested more complex training procedures but found no improvement in performance. In addition, each step uses the variables defined for that stage, for example, the super-resolution model only uses $Z$ as input and not also $X$. This simplification approximates the full distribution $p^{\iota}_{Y|X}$ accurately if $Z$ contains all the relevant information about the target $Y$ or, in other words, the map from $Z$ to $Y$ does not depend on $X$ anymore. This approximation is formalized through conditional independence and checked heuristically in App. B.1. Empirically, we found no (e.g., precipitation) or at most 2 % (e.g., temperature) decrease in performance due to this approximation.
Figure 2: Downscaling via coarse correction for EnScale. We approximate the conditional distribution of RCM data $Y$ given GCM data $X$ with a two-step approach. In the second row, $Z$ represents RCM data manually coarsened through average pooling. The map learning the conditional $p_{Z|X}$ is called the coarse model, and the map for the conditional $p_{Y|Z}$ the super-resolution model. All $X,Z,Y$ include multiple climate variables and grid points, and represent daily data.
2.3 Modeling temporal consistency
So far, we predicted each RCM day independently, ignoring the temporal context. However, our architecture can naturally be extended to generating temporally consistent samples. For time series generation, we aim to learn the distribution of the RCM data $Y_{t}$ for a given day $t$, given the GCM input $X_{t}$ on the same day and the RCM data $Y_{t-1}$ from the previous day. We restrict temporal dependence to one-day lags, similar to the Markov assumption in ML-based weather forecasting [Lam2023LearningForecasting, Price2025ProbabilisticLearning]. More long-term memory effects are potentially still present in the GCM input.
Conditioning on $Y_{t-1}$ directly is challenging because $Y_{t-1}$ is high-dimensional. Thus, we approximate the temporal dependency on the lower-dimensional scale. For this, we train a second temporal coarse model to model $Z_{t}$ given $Z_{t-1}$ and $X_{t}$. This models autocorrelation on large spatial scales, e.g. in pressure systems. During training, we use the average pooled version of the RCM data $Y_{t-1}$ as predictors for $Z_{t}$. During inference, we generate a time series with an “autoregressive rollout”: Given a sample $\hat{Z}_{t-1}$ for time step t1t-1, we predict $\hat{Z}_{t}$ with the temporal coarse model using $\hat{Z}_{t-1}$ and $X_{t}$. This is repeated iteratively for future time steps. For the initial time step $t=1$, we use one sample from the marginal coarse model, only based on $X_{1}$.
Figure 3: Time series generation with temporal consistency in EnScale-t using an autoregressive roll-out. At each time step $t$, the model predicts coarsened RCM data $\hat{Z}_{t}$ using the sample from the previous day, $\hat{Z}_{t-1}$, and the GCM data from the same day, $X_{t}$. The super-resolution model is applied to independently to each $\hat{Z}_{t}$ to generate the final high-resolution time series. For simplicity, this figure presents only one variable (precipitation), but the approach uses all variables jointly.
The resulting time series $\hat{Z}_{t}$ is then passed through the super-resolution model, without temporal dependency, to obtain a time series for $Y_{t}$ (Fig. 3). This approximation neglects temporal correlations on a local scale. We check the effect of this approximation in App. B.2, finding that the approximation does not degrade performance at all (e.g. precipitation) or by at most 2 % (e.g. surface wind).
Our approach to model temporally consistent time series approximates the distribution $p^{\iota}_{Y_{t}|X_{t},Y_{t-1}}$ with a few conditional independence statements, see App. B.2 for details.
2.4 Generative model architecture
We extend model architectures to our stochastic and high-dimensional setting by combining dense and sparse architectures. The coarse model uses a dense multilayer perceptron (MLP), following [Shen2024Engression:Regression]. However, dense layers are very memory-intensive for high-dimensional data. Thus, we design a sparse architecture for the high-dimensional super-resolution steps.
Firstly, the complex and high-dimensional super-resolution task of learning $p^{\iota}_{Y|Z}$ is split into four simpler steps, as in [Shen2025ReverseDistributions]. In particular, each step increases resolution by a factor of two in both spatial dimensions, i.e., we model $p^{\iota}_{\tilde{Z}|Z^{\prime}}$, where $Z^{\prime},\tilde{Z}$ are average-pooled versions of the target $Y$ with dimensions $2^{k}\times 2^{k}$ and $2^{k+1}\times 2^{k+1}$, respectively. The first step uses $k=3$, corresponding to the dimension of $Z$ ($8\times 8$), and $k$ increases by 1 for each step, so that the last step uses $k=6$, as $Y$ has dimension $128\times 128$.
Secondly, each conditional distribution $p^{\iota}_{\tilde{Z}|Z^{\prime}}$ is modeled with sparse local layers. The value of $\tilde{Z}$ in one location is approximated to only depend on the values in a local neighborhood grid cells in $Z^{\prime}$, with learnable, location-specific weights. This design allows the model to capture local characteristics such as topography or land-sea contrast. At the same time, by restricting interactions to local neighborhoods and by sharing non-linear transformations between locations, we keep the model size manageable even for high resolutions. Our approach combines flexibility of dense connections with efficiency of convolutional architectures.
For our generative modeling task, the ML model must generate multiple outputs for one single input $X$. To achieve this, we integrate Gaussian noise into our model as a source of stochasticity. We do this separately for each subpart of our model (coarse model and each component of the super-resolution model). For the dense MLP architectures, the noise is concatenated to hidden units in each hidden layer. This facilitates learning complex representations of the random inputs and yields spatially correlated variability in the samples, in line with independently developed approaches such as [Alet2025SkillfulMarginals], [Lang2024AIFS-CRPS:Score]. For the sparse layers, the stochastic super-resolution proceeds in two stages (see Fig. 4). A first deterministic upsampling step is done separately for each target variable, and interpolates each high-resolution pixel in $\tilde{Z}$ linearly from its nearest low-resolution neighbors in $Z^{\prime}$ with learnable location-specific weights. The second step performs a stochastic refinement. The intermediate upsampled variables are stacked and concatenated with several channels of Gaussian noise in each pixel; this concatenation is used as input to another local aggregation. Each pixel again has its distinct set of linear weights, applied to inputs from a local neighborhood, using both the upsampled variables as well as the noise channels. This step enables generating spatially correlated variability and capturing multivariate dependencies. Finally, a shared MLP is used for all pixels (weight-sharing across space), where each pixel is processed independently of others, but all dimensions are considered jointly. More details on all model architectures are given in App. B.4.
Figure 4: Sparse local stochastic layers from EnScale’s super-resolution model. As an example, we demonstrate modeling the distribution in an example pixel of interest (top left corner, orange); the same procedure is applied to all pixels. For each intermediate map (small arrows), light blue shaded pixels serve as inputs and orange pixels as the targets. First, a deterministic upsampling step processes each target variable (tas, pr, sfcWind, rsds) separately, linearly interpolating nearest pixels in the low-resolution input with learnable weights. Second, the intermediate upsampled variables are stacked and concatenated with noise channels. Next, each pixel again interpolates linearly from its nearest neighbors with learnable location-specific weights (this time using all variables and the noise as inputs). Finally, an MLP is applied to each pixel independently, using the same weights at each location.
Our choice of loss function ensures that the model indeed learns the correct conditional distributions, as we describe next.
2.5 Generative Models with Proper Scoring Rules
2.5.1 Connection to forecasts and forecast evaluation
Our method, along with our benchmarks and evaluation techniques, is inspired by the correspondence between generative modeling and forecast evaluation. Consider the example of weather forecasts. Given the state of the atmosphere at a certain time, weather forecasts aim to predict the weather a few hours ahead. However, rather than issuing a single best estimate, they also aim to quantify uncertainties by predicting the full probability distribution of possible future weather for the next time step [Gneiting2005WeatherMethods]. Typically, this is done by running an ensemble of simulations from a numerical weather model. Afterwards, a post-processing step corrects biases and variance discrepancies of the ensemble [Vannitsem2021StatisticalWorld]. This post-processing is trained based on previous forecasts and corresponding observations. The goal is to produce a probability distribution that should be consistent with the target, i.e., the observed data.
A key challenge in evaluating such weather forecasts is that, at each time step, only a single ground truth from observations is available. Thus, tools are needed to compare a single observed value with the full predicted distribution. Averaging over multiple forecast-observation pairs then allows evaluating the quality of the distributional forecasts [Gneiting2007ProbabilisticSharpness].
This setup is analogous to generative modeling: for each GCM input, the generative ML model outputs an ensemble of predictions (multiple samples). The model’s samples serve as an approximation to the underlying (continuous) probability distribution of possible outcomes. We wish to evaluate these predictions relative to the target data from the RCM, where for each RCM we only have a single realization available. This comparison between forecast evaluation and generative modeling is crucial both during training of the generative model, as well as in its evaluation.
2.5.2 Energy score loss
For simplicity, we first consider the unconditional generation of a target distribution $p_{Y}$ and neglect GCM predictors $X$. We denote the distribution from the ML model by $q_{Y}$. Recall that we do not have explicit access to the ML’s output distributions, but instead have samples for $p_{Y}$ and can generate samples for $q_{Y}$ via the generative ML model. We denote true data by $Y$ and two exemplary predictions from the model by $\hat{Y},\hat{Y}^{\prime}$.
[…]
The total loss then presents a trade-off between prediction error (first term) and the scaled variability of the samples (second term)
$\text{Loss}_{\text{uncond}}=E_{Y\sim p_{Y},\hat{Y}\sim q_{Y}}(\lVert\hat{Y}-Y\rVert)-\frac{1}{2}E_{\hat{Y},\hat{Y}^{\prime}\sim q_{Y}}(\lVert\hat{Y}-\hat{Y}^{\prime}\rVert).$ (1)
Here, the subscript “uncond” refers to the fact that the loss does not take into account any conditioning yet. The norms in this score are Euclidean norms that can be applied directly to high-dimensional targets $Y$. Both $\hat{Y}$ and $Y$ typically are full spatial fields for multiple climate variables at once. When combining multiple variables with different physical units, a data transformation for standardization is crucial, as we introduce in Sec. 3.2.
Our loss can uniquely identify the correct distribution: $\text{Loss}_{\text{uncond}}$ is minimal if and only if the modeled distribution $q_{Y}$ follow the target distribution $p_{Y}$ ($q_{Y}=p_{Y}$). More precisely, it holds that, for all $p_{Y},q_{Y}$
$E_{Y\sim p_{Y},\hat{Y}\sim q_{Y}}(\lVert Y-\hat{Y}\rVert)-\frac{1}{2}E_{\hat{Y},\hat{Y}^{\prime}\sim q_{Y}}(\lVert\hat{Y}-\hat{Y}^{\prime}\rVert)$
$\geq E_{Y\sim p_{Y},\hat{Y}\sim p_{Y}}(\lVert Y-\hat{Y}\rVert)-\frac{1}{2}E_{\hat{Y},\hat{Y}^{\prime}\sim p_{Y}}(\lVert\hat{Y}-\hat{Y}^{\prime}\rVert),$
and equality holds if and only if $p_{Y}=q_{Y}$ [Rizzo2016EnergyDistance, SzekelyE-Statistics:Distances]. This implies that minimizing this loss ensures that the generated samples follow the target distribution.
If $p_{Y}=q_{Y}$, both terms of the expected energy score are indeed equal, i.e.,
$E_{Y\sim p_{Y},\hat{Y}\sim q_{Y}}(\lVert\hat{Y}-Y\rVert)=E_{\hat{Y},\hat{Y}^{\prime}\sim q_{Y}}(\lVert\hat{Y}-\hat{Y}^{\prime}\rVert).$
However, the contrary is not true, i.e., equal scores do not imply that the score is minimal. In other words, equality is a necessary but not sufficient criterion for reaching the minimum.
This loss is related to the energy score in forecast evaluation. For a single target data point $y$ and a predicted distribution $q_{Y}$, the energy score is defined as
$S(q_{Y},y)\coloneqq E_{\hat{Y}\sim q_{Y}}(\lVert y-\hat{Y}\rVert)-\frac{1}{2}E_{\hat{Y},\hat{Y}^{\prime}\sim q_{Y}}(\lVert\hat{Y}-\hat{Y}^{\prime}\rVert).$
This is a multivariate generalisation of the Continuous Ranked Probability Score (CRPS) [Gneiting2007StrictlyEstimation, SzekelyE-Statistics:Distances] frequently used in numerical weather prediction. For one-dimensional $Y$, the energy score and CRPS are equivalent. The CRPS is typically applied per variable and per grid point, whereas we use the energy score jointly over all locations (and, unless stated otherwise, also across all target variables), capturing dependencies between them.
The above loss function $\text{Loss}_{\text{uncond}}$ corresponds to the expected energy score over multiple target data points, for example several days in the RCM: $\text{Loss}_{\text{uncond}}=E_{Y\sim p_{Y}}(S(q_{Y},Y))$.
2.5.3 Conditional case
Our generative model should learn to sample from the RCM’s conditional distribution $p^{\iota}_{Y|X=x}$ for all GCM inputs $x$ and for each GCM-RCM pair $\iota$. Thus, we need to adjust our loss function to this conditional case. For simplicity, we drop the superscript $\iota$ in the following description. We define the loss for inputs $X$ and output $Y$, but apply the equivalent loss function to the individual steps in our pipeline (see Sec. 2.4). The generative model, conditioned on $X$, samples from the conditional distribution $q_{Y|X}$.
The objective is to ensure that the learned conditionals match with those from the RCM, i.e., $q_{Y|X=x}=p_{Y|X=x}$ for all $x$ and for each RCM. However, for each GCM input $x$ and for each RCM, we only have one corresponding RCM realization $Y\sim p_{Y|X=x}$. Thus, we draw multiple GCM inputs $X$ from the GCM distribution $p_{X}$ and average over those in the loss function:
$\text{Loss}_{\text{cond}}=\underbrace{E_{X\sim p_{X}}\Bigg(E_{Y\sim p_{Y|X},\hat{Y}\sim q_{Y|X}}(\lVert Y-\hat{Y}\rVert)\Bigg)}_{\text{ES}_{\text{pred}}}-\frac{1}{2}\underbrace{E_{X\sim p_{X}}\Bigg(E_{\hat{Y},\hat{Y}^{\prime}\sim q_{Y|X}}(\lVert\hat{Y}-\hat{Y}^{\prime}\rVert)}_{\text{ES}_{\text{var}}}\Bigg)$ (2)
We call this the EnScale-loss. We call the prediction error (first term) $\text{ES}_{\text{pred}}$ and the variability (second term) $\text{ES}_{\text{var}}$. This loss is minimized if and only if, for each predictor $x$, the predicted conditional distributions $q_{Y|X=x}$ indeed match the target conditionals $p_{Y|X=x}$. More formally, the score is minimized if and only if $q_{Y|X=x}=p_{Y|X=x}$ $p_{X}$-almost everywhere [Pacchiardi2024ProbabilisticMinimization]. As for the unconditional loss in Eq. (1), if $p_{Y|X=x}=q_{Y|X=x}$ for each $x$, it holds that $\text{ES}_{\text{pred}}=\text{ES}_{\text{var}}$. During training, we use minibatching. We describe the standard batch-wise approximation of the loss in App. B.3, where we use two samples by the generative model for each input $x$. Each batch contains data from multiple GCM-RCM pairs, such that the loss is averaged across pairs.
We add additional predictors for seasonality information. The day of the year is passed as an integer together with its sine and cosine to facilitate learning seasonal information. We do not input the year explicitly in order to allow for extrapolation to years that have not been seen in the training data.
3 Data
3.1 Data preprocessing
RCM GCM
ALADIN63 CNRM-CM5
ALADIN63 MPI-ESM-LR
CCLM4-8-17 CNRM-CM5
CCLM4-8-17 MIROC5
CCLM4-8-17 MPI-ESM-LR
REMO2015 MIROC5
RegCM4-6 CNRM-CM5
RegCM4-6 MPI-ESM-LR
Table 1: Regional climate model and global circulation model pairs (EURO-CORDEX) used in our study.
We use data from eight GCM-RCM pairs from the EURO-CORDEX framework [Jacob2014EURO-CORDEX:Research], see Table 1. We consider daily 2m temperature (tas), total precipitation (pr), surface wind (sfcWind) and surface downwelling shortwave (solar) radiation (rsds) from both driving GCMs and corresponding downscaled RCMs. In addition, we add sea level pressure (psl) from the GCMs as a predictor because it provides valuable information about the large scale weather pattern. The selection of climate models was based on the availability of daily surface wind data, which not all models publish. We chose a set of driving GCMs for which multiple corresponding downscaled datasets from different RCMs were available. We use daily data for the years 1971-2099, from both the historical and the RCP8.5 experiment.
3.1.1 Targets
The targets are the RCM data in Central Europe for all four variables mentioned above. The domain spans $0^{\circ}E$ to $18.9^{\circ}E$ and $42^{\circ}N$ to $54.8^{\circ}N$ and includes land as well as sea areas and the Alps. For most of the RCMs output, the data are provided on the EUR-11 grid. Otherwise, we re-gridded them with first-order conservative remapping.
Different RCMs write data in different units. We rescaled temperature to Celsius, precipitation to mm/day, solar radiation to $\mathrm{W}/\mathrm{m}^{2}$ and wind to m/s.
When both the driving GCM and the corresponding RCM dataset include the 29th of February in leap years, we left the data as is. However, when the GCM uses a no-leap calendar with 365 days in each year and the RCM has a calendar including leap days, we removed all 29th of February in the RCM data. After preprocessing, driving GCM and corresponding RCM dataset always have the same length.
3.1.2 Predictors
The inputs for the ML model are the GCM data from the driving GCM. The GCMs were regridded via second-order conservative remapping to a regular 2.5 degree grid. We select a domain which is larger than the target domain and covers Europe, spanning $40^{\circ}W$ to $50^{\circ}E$ in longitude and $25^{\circ}N$ to $75^{\circ}N$ in latitude.
We tested the addition of further variables as predictors, e.g. humidity, wind and geopotential height at 500 and 850 hPa, but found no improvement in performance on the test loss.
3.1.3 Train-test split
The total runs span the years 1971-2099, and we use 1971-2029 and 2040-2089 as training data. A random subset of 10 % of the training data is taken as the validation set. The two test data sets are 2030-2039 and 2090-2099. The first test data set represents the “interpolation” test case, whereas the second is the “extrapolation” case with test data outside the time period covered in the training data. This yields approximately 320k data points for training and around 30k points for each test set.
3.2 Data transformation for training
To combine multiple variables in a loss function for model training, the data must be transformed to avoid issues arising from different scales or units. For the target data from the RCM, we transform each variable and location to a normal distribution.
First, we approximate the empirical distribution function (ECDF), $\hat{F}$, on a subsample of 10 % of the training data. The RCM’s target data from all GCM-RCM pairs are pooled, but we fit the ECDF separately for each variable and location. Applying the ECDF to the training data, again in a univariate manner, yields probabilities $\hat{F}(Y)$ that are uniformly distributed on $(0,1)$. In the case of precipitation, however, the ECDF introduces a point mass at zero rainfall and the uniformity only holds above this threshold. Second, we map to a normal distribution by computing $\Phi^{-1}(\hat{F}(Y))$, where $\Phi$ is the CDF of the standard normal distribution. This maps the probabilities $\hat{F}(Y)$ from $(0,1)$ to $\mathbb{R}$ and implies that the target values used in ML training lie in $\mathbb{R}$. This transformation ensures that all locations and variables follow the same distribution, allowing joint application of the loss to all variables.
[…]
For the input data, the precise transformation is less crucial, so we apply a simpler and computationally more efficient normalization: each variable is standardized by subtracting the mean and standard deviation. Here, mean and SD were computed using the entire spatial field and training data, again from all GCMs.
4 Methods for evaluation
4.1 Benchmarks
4.1.1 Deterministic neural network
The most simple deterministic benchmark is a deterministic neural network (NN-det in the following) with an MSE loss. As input, we use five GCM variables, the four target variables plus sea level pressure. For simplicity, we have separate ML models for each target variable. More details are given in App. B.6.
4.1.2 EasyUQ
EasyUQ is a tool to convert predictions from a deterministic statistical model into probabilistic ones [Walz2024EasyOutput]. We use it to obtain predicted probability distributions from the above simple NN-det (Sec. 4.1.1). We introduce EasyUQ because, based on statistical theory, it is expected to issue good probabilistic predictions for each grid point and variable separately, as a baseline comparison for the performance of EnScale on at the level of individual locations. However, EasyUQ breaks spatial and inter-variable dependencies in the output.
EasyUQ it is applied as a post-processing step after training the NN-det, independently for each GCM-RCM pair. The NN-det’s predictions, denoted by $\hat{\mu}$, approximate the conditional mean $\mu\coloneqq E[Y|X]$. Given training data consisting of pairs of NN-det’s predictions $\hat{\mu}$ and the RCM target data $Y$, EasyUQ quantifies the uncertainty in the NN-det’s predictions. It returns full probability distributions. More details on theoretical properties and the implementation are given in App. B.6.
4.1.3 Analogues
Analogue-based downscaling [Boe2023AValue] is a statistical-dynamical downscaling method which resamples the existing RCM data. It is based on a nearest neighbor search in the GCM data, and stochastic downscaling is approximated by selecting five similar GCM days and using their corresponding RCM days as probabilistic predictions. The main strength of the analogue approach is that, by construction, it yields samples with the correct spatial pattern, as they are all drawn from the existing RCM data. However, it only reproduces the training data and can never create truly new random fields, and the predicted conditional distributions are only discrete. Finally, due to a lack of perfect “near neighbors”, the analogue benchmark is expected to have a larger prediction error between target RCM data and predicted samples than the true RCM’s conditional distribution. Details on our algorithm and potential limitations are given in Sec. B.7.
4.1.4 Generative Adversarial Network
Generative Adversarial Networks (GANs) have successfully been used to downscale precipitation fields [Harris2022AForecasts, Leinonen2020, PriceStephanRasp2022IncreasingModels]. As an additional benchmark, we consider the approach presented in [Harris2022AForecasts], which is a Wasserstein GAN with gradient norm penalty (WGAN-GP). In the generator objective function, it includes an additional content loss term, more specifically, the mean-squared error of the mean prediction (over eight generated samples per input field) and the ground truth high-resolution output. We adjust [Harris2022AForecasts] [^1] Code: https://github.com/ljharris23/public-downscaling-cgan to our multivariate setting. As for EnScale, we use five input fields (tas, pr, sfcWind, rsds, psl) as the conditioning inputs to the WGAN-GP. Additionally, we include the one-hot-encoded index of the considered GCM-RCM pair and concatenate it to the five input fields. The outputs are the four target fields (tas, pr, sfcWind, rsds). Similar to [Harris2022AForecasts], we scale the precipitation and surface wind values logarithmically with $\log(1+x)$. In addition, we use min-max normalization to ensure that all variables have values on similar scales. For implementation reasons, we confine the input domain of the GCM inputs to $11.25^{\circ}E$ to $26.25^{\circ}E$ and $31.25^{\circ}N$ to $68.75^{\circ}N$, i.e., a smaller extent than covered in EnScale. We use the same hyperparameters as in [Harris2022AForecasts] with a larger batch size of 16 and train until convergence, reached after 2.4 million training samples.
4.1.5 Corrective Diffusion Model
As another state-of-the-art ML downscaling approach, we consider CorrDiff, which was shown to perform skillfully in multivariate downscaling [Mardani2025ResidualDownscaling]. The model consists of a UNet to predict the (deterministic) conditional mean $E[Y|X]$ and a diffusion model to predict the residual variation around the mean. We adopt the CorrDiff implementation provided in NVIDIA’s PhysicsNeMo library [^2] Code: https://github.com/NVIDIA/physicsnemo/tree/main/examples/weather/corrdiff for the same setting as in the previous GAN experiment, i.e., five input fields concatenated with the one-hot-econded GCM-RCM pair indices, four output fields, and the confined GCM-input-domain.
As in [Mardani2025ResidualDownscaling], we standardize all variables to z-scores and use the same hyperparameter settings. In particular, we use a base embedding size of 128 for the UNet, channels are multiplied with the factors $[1,2,2,2,2]$, and the attention resolution is defined as 8 (as our input is of size $128\times 128$). We use a smaller batch size of 32, and train until convergence: first, the UNet for 4 million training samples and, afterwards, the diffusion model for 26 million training samples.
4.2 Validation
To design our validation metrics, we again leverage the connection between generative models and forecast evaluation (Sec. 2.5.1). Two important properties of the generated samples by the ML model are calibration and sharpness [Gneiting2007ProbabilisticSharpness].
Calibration is also called “reliability” as it checks whether the model’s samples have the right “spread” and correctly capture all possible outcomes compared to the target RCM data. In other words, a well-calibrated model should not systematically over- or under-represent typical or extreme values. However, calibrated samples are not necessarily informative: for example, if the ML model ignored the GCM input $X$ and generated samples from the climatology or from the marginal distribution across all days of the year, independently of $X$, the samples would be calibrated. But the samples are not very useful, as they carry little information about the particular day of interest.
Samples with small variance are called sharp. Sharpness does not imply that the samples are correct: a narrow range that is always far away from the truth would also be sharp. Calibrated samples balance prediction error and variability and often achieve $\text{ES}_{\text{pred}}=\text{ES}_{\text{var}}$, but only samples that are also sharp (small $\text{ES}_{\text{var}}$) are optimal (see our explanation of the loss, Sec. 2.5.2).
We assess calibration and sharpness for the samples of our model in general and with a particular focus on extremes. In addition, we consider and temporal structure, as well as variable dependencies.
4.2.1 Energy score and CRPS
To evaluate the ML model, we use proper scoring rules, in particular the CRPS and the energy score. These assess calibration and sharpness jointly.
We consider the multivariate EnScale-loss for the full spatial field for each variable separately (see Sec. 2.5.3). In addition, we calculate CRPS values, which coincide with the energy score in the univariate case. We calculate the CRPS for each variable and in each location and then average it across the spatial field.
4.2.2 Spatial structure
To check the spatial structure, we combine the locations-wise CRPS calculation with a pre-pooling operation. We first apply max-pooling over patches of the data and then compute CRPS values for the pooled patches. Results are shown for a kernel size of 10. The conclusions are consistent across different kernel sizes.
In addition, we compute power spectral densities (PSD). We follow [Harris2022AForecasts] and consider the radially-average PSD, calculated with pysteps. Following [Harris2022AForecasts], we compare the sampled PSDs with the RCM’s PSDs via the log spectral distance, called RALSD:
$\text{RALSD}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\Bigl(10\text{log}_{10}\frac{\overline{P}_{\text{RCM, i}}}{\overline{P}_{\text{gen, i}}}\Bigr)^{2}}$ (3)
Here, $\overline{P}_{\text{RCM, i}}$ and $\overline{P}_{\text{gen, i}}$ are the power spectra for the RCM data and one generated field at one wavelength $i$, where the 2D spectra were averaged radially. $N$ is the number of wavelengths considered. The scaling by factor 10 is chosen such that the resulting units are in decibel. To arrive at RALSD, we compute PSDs separately for each test day and compute the RALSD on daily PSDs as in Eq. (3). Finally, we average over all days in the test set, as in [Harris2022AForecasts].
However, the PSDs $\overline{P}_{\text{gen, i}}$ for different random samples can vary substantially even on the same day, due to the large variability in the conditional $p_{Y|X}$. Thus, RALSD is also sensitive to general mismatches between samples and target, and even a method with correct spatial structure might score poorly if its realizations deviate from the RCM target. Hence, we consider averaging the radially-averaged PSDs across all days before compute the log spectral density. This version emphasizes systematic errors in spatial structure, and we refer to this metric as RALSD-avg.
4.2.3 Calibration
We use rank histograms (RHs) to evaluate the calibration of our model’s samples. RHs consider ranks of the target RCM data $y$ within an ordered set of samples, e.g. $y_{1}<...<y_{9}$. These ranks are visualized in a histogram. If the samples are calibrated, all ranks from 1 to 10 are equally likely, implying a flat RH. This simultaneously checks if the samples are unbiased compared to the truth and if they have the right variability, i.e., are neither over- nor underconfident. Deviations from flatness indicate which type of violation occurs: for example, a U-shaped histogram suggests overconfident samples (too little variance), while a dome-shaped histogram indicates underconfident samples (too much variance).
RHs can only consider one-dimensional targets and samples. Thus, we reduce the high-dimensional spatial field to a one-dimensional quantity by taking the mean or maximum before plotting. Additionally, we compute rank histograms at each location and quantify miscalibration using the miscalibration measure (MCB). The MCB is calculated as the sum of the absolute deviations between the generated RH and a uniform (flat) rank histogram. Thus, an MCB of zero indicates a perfectly calibrated model, and larger values indicate increasing miscalibration. By averaging the MCB across all locations, we obtain a general measure of calibration quality. However, the MCB can not assess which type of violation occurs, e.g. whether the samples are over- or underdispersed.
4.2.4 Extremes
We consider extreme quantiles across all days in a season. For example, we compute the 99 % - quantile for all samples on summer days in the test set. These quantiles are computed in each location separately. Afterwards, the absolute errors of the sampled vs. true RCM’s quantiles are averaged across all locations.
Additionally, we analyze behaviour in the extremes of the predicted conditional distribution with a miscalibration measure, focussing on the calibration of the most extreme samples. We investigate calibration of high extremes in a single location as follows: for each day, we consider 9 samples of the generative models, $y_{1},...,y_{9}$, select the highest value, say $y_{9}$, and compare it to the target RCM data $y$. If the samples capture extremes correctly, the target RCM data $y$ should lie above $y_{9}$ in 10 % of the days. Hence, we calculate the ratio of occurrences of the case $y_{9}<y$ (out of all days) and then consider the absolute deviation from this ratio to 0.1. This is a miscalibration measure similar to the general MCB metric from Sec. 4.2.3, but now only considers the last bin of the RH. We average the extremal MCB scores across all locations, and proceed likewise for the low extremes.
4.2.5 Temporal structure
We evaluate the temporal structure of the samples by computing auto-correlations. For each variable and each location, we calculate temporal lag-1 autocorrelation for the generated time series as well as the RCM target, and report absolute errors of the samples.
4.2.6 Variable dependencies
We evaluate pairwise dependencies between several pairs of the four target variables. Firstly, we consider Pearson correlation coefficients in each location for both the RCM as well as the samples. We report absolute errors of the correlation coefficients, averaged across all locations.
In addition, we consider a novel metric of calibration: we compute RHs and MCB metrics (see Sec. 4.2.3) for the difference of two variables in the samples. However, as directly comparing variables with different physical units is not meaningful, we first apply a data transformation to standardize them. Specifically, for each variable and each dataset (RCM or sampled), we transform the data to a uniform distribution. This ensures that the marginal distributions for each variable are identical, eliminating differences to units. In addition, it removes mismatches between RCM and sampled data in individual variable distributions, such that the remaining discrepancies arise only from different dependence structures between the variables. After transformation, we calculate the differences between two variables. We compute RHs and MCBs for each grid point separately and then averaged across locations.
4.3 Contributions to variability in EnScale
We aim to better understand the variability EnScale’s samples and present decompositions of its conditional variance $\operatorname{Var}(Y|X)$. As discussed (Sec. 2.1), there is a wide range of possible realizations of high-resolution fields given a particular GCM input. EnScale allows us to quantify this variability. In this section, all variances and expected values refer to those of EnScale’s samples, assuming already that EnScale emulates the true RCM’s distributions accurately. In addition, $\operatorname{Var}$ and $E$ denote quantities for the pooled distribution across multiple GCM-RCM pairs, $\iota$ denotes the index for a GCM-RCM pair, and $\operatorname{Var}^{\iota}$ resp. $E^{\iota}$ refer to distributions by an individual pair.
4.3.1 Coarse-scale vs. local-scale variance
We aim to decompose the internal variability of EnScale’s full GCM-RCM path into the variability on coarse and local scales. For this, we split the variance of EnScale, $\operatorname{Var}^{\iota}(Y|X)$, into contributions from (i) variance of the coarse model, $\operatorname{Var}^{\iota}(Z|X)$, which we call coarse-scale variability, and (ii) variance of the super-resolution model, $\operatorname{Var}^{\iota}(Y|Z)$, which we call local-scale variability (see Sec. 2.2). As $Z$ and $Y$ have different dimensions and physical meanings, $\operatorname{Var}^{\iota}(Z|X)$ and $\operatorname{Var}^{\iota}(Y|Z)$ are not directly comparable. Instead, we can formally decompose the conditional variance $\operatorname{Var}^{\iota}(Y|X)$ as follows:
$\operatorname{Var}^{\iota}(Y|X)=\underbrace{E^{\iota}[\operatorname{Var}^{\iota}(Y|Z)|X]}_{\text{local-scale variance}}+\underbrace{\operatorname{Var}^{\iota}(E^{\iota}[Y|Z]|X)}_{\text{coarse-scale variance}},$ (4)
where the first term is the variance of $Y|Z$, averaged over all possible $Z$ (conditional on $X$). The second term describes the variance of the conditional mean $E^{\iota}[Y|Z]$, again for different possible $Z$. This is conceptually very similar to $\operatorname{Var}^{\iota}(Z|X)$, except it is “evaluated in YY”. A proof is given in App. B.8. This decomposition holds separately in each grid point and for each variable. We average each of the terms over all grid points as well as over all $X$ in the test set.
We calculate the quantities above as follows: for a given GCM input $X$, we generate 9 realizations of $Z$ with the coarse model. For each $Z$, we again generate 9 realizations of $Y$ with the super-resolution model, and compute $\operatorname{Var}^{\iota}(Y|Z)$ and $\operatorname{Var}^{\iota}(E^{\iota}[Y|Z]|X)$ for each grid point. Afterwards, we average across $X$, grid points and GCM-RCM pairs. To restrict memory usage, the calculation is only done for the first year in the test set.
4.3.2 Internal variability vs. model differences
EnScale allows us to compare the magnitude of model differences between RCMs with the intrinsic variability in individual RCMs. As an example, consider a fixed GCM run that is dynamically downscaled by two RCMs. Assume we train EnScale to emulate both RCMs, such that it can generate samples from each of them, yielding a pooled dataset. Using these samples, we quantify contributions to the variance in this pooled data, $\operatorname{Var}(Y|X)$, from (i) internal variability, present in each RCM, and (ii) structural differences between the two RCM.
Formally, let the first RCM correspond to index $\iota=1$ and the second RCM to $\iota=2$. For example, $\operatorname{Var}^{\iota=1}(Y|X)$ denotes the variance of the first RCM. We decompose $\operatorname{Var}(Y|X)$ as follows:
$\operatorname{Var}(Y|X)=$ $\underbrace{\frac{1}{2}\big(\operatorname{Var}^{\iota=1}(Y|X)+\operatorname{Var}^{\iota=2}(Y|X)}_{{\text{Internal variability in individual RCMs}}}\big)$ (5)
$+\underbrace{\frac{1}{2}\big((E^{\iota=1}[Y|X]-E[Y|X])^{2}+(E^{\iota=2}[Y|X]-E[Y|X])^{2}}_{\text{Model differences between RCMs}}\big).$ (6)
Thus, the total variance can be written as an average of the variances of the individual RCMs (internal variability) plus the squared deviations of the means of each RCM from the mean of the pooled data (model differences). This decomposition analogous to classical ANOVA. For a proof, see [Casella2024StatisticalInference]. As above, we consider this decomposition separately per grid point and per variable, and average over grid points and predictors $X$.
5 Evaluating the performance of EnScale
5.1 Comparative overview of method performances
Figure 5: Summary of performance of EnScale compared to the benchmarks in several selected categories, shown for the interpolation test period (2030-39). Energy score (see Sec. 5.4.1), Calibration (see Sec. 5.4.3), Spatial structure (see Sec. 5.4.2), Temporal structure (see Sec. 5.4.5), Extremes (see Sec. 5.4.4), Multivariate dependencies (Sec. 5.5). The chosen metrics for the categories are outlined in more detail in the main text. All metrics are normalized such that EnScale attains a score of 1 in the respective category and all other scores are expressed relative to EnScale. In all categories, lower values indicate better performance. As calibration for deterministic models is not meaningful, we do not show it for NN-det.
Figure 6: Time series examples. Solid lines show the RCM time series (green) and a single randomly chosen realization of EnScale (blue), EnScale-t (yellow), and CorrDiff (orange). Shaded areas indicate the minimum and maximum from 9 samples of each model. The figure shows spatially averaged temperature for the first year in the interpolation test set (2030) and for the first GCM-RCM pair (CNRM-CM5 and ALADIN63).
Fig. 5 summarizes the performance of EnScale compared to the benchmarks for several important categories. “Energy score” refers to the spatially multivariate energy score, “Spatial structure” is calculated from the max-pooled CRPS and the radially-average log spectral density (RALSD and RALSD-avg), “Temporal structure” refers to the autocorrelation functions, and “Extremes” contains the marginal quantiles and quantile losses. To evaluate variable dependencies, we compute correlations for multiple variable pairs and assess calibration of their differences. The results are then averaged and summarized as “correlations” and “calibration”. All metrics are analyzed in more detail in the respective sections below. Unless otherwise stated, all metrics are shown on the interpolation test set (2030-2039) and averaged across all GCM-RCM pairs.
We find that EnScale and EnScale-t reach very good scores for many categories and variables in comparison to the benchmarks. It particularly stands out with respect to calibration, extremes as well as the multivariate dependencies where it outperforms benchmarks by a large margin.
CorrDiff yields mixed results: despite good energy score and realistic spatial structure, its performance for calibration, extremes and multivariate dependencies is poor. We find that these deficiencies are a consequence of the sampling scheme in CorrDiff, which is intended to provide ensemble members. The current implementation yields highly autocorrelated time series and restricts the diversity of samples. Thus, we adapted the sampling to allow a broader exploration of the space of possible high-resolution fields. This new version, which we call CorrDiff-s, substantially improves on CorrDiff’s performance in most metrics, as shown in App. A.2.
The analogues benchmark performs well in many metrics as is to be expected. This benchmark resamples existing fields. For this reason, it perfectly reproduces the spatial characteristics and variable correlations of the RCM, thus reaching excellent scores in these categories. The GAN is inferior compared to the other ML-approaches. EasyUQ performs well on metrics that consider each location separately, but leads to the worst performance regarding calibration and spatial structure.
Fidelity regarding the temporal structure is of particular relevance and the extension to EnScale-t exhibits the best scores of all stochastic methods in that regard. This is because EnScale-t is the only method that explicitly models temporal structure. All other methods downscale each day independently of previous days, with CorrDiff being one exception due to shared random initializations between different time steps. Generated time series also show that EnScale-t and CorrDiff produce smoother, and thus more realistic, trajectories than EnScale (see Fig. 6).
We compare EnScale (EnScale-t) to EnScale-no-label (EnScale-t-no-label) which exclude the one-hot-encoded index of the GCM and RCM in the super-resolution model, thus learning a joint super-resolution map for all pairs (see Fig. A.1). EnScale yields small improvements to EnScale-no-label. Overall, results are very similar, again underlining that the super-resolution map behaves similarly for multiple RCMs.
In addition, we repeat all metric calculations on the extrapolation test period (2090-2099) and show the result in App. A.4. As this is only a mild extrapolation case (as training data is until 2089), scores are very similar to the interpolation period and the ordering of the methods stays consistent.
5.2 Computational cost
Method Training Inference
EnScale 24 h 1.75 h
CorrDiff 216 h 36 h
Table 2: Training and inference times for EnScale and CorrDiff. All runtimes were checked on a single NVIDIA A100 GPU (80GB) and exclude data loading. Inference times refer to producing a high-resolution ensemble of 100 years for a single GCM-RCM pair with 10 samples per day.
EnScale aims to be a computationally lightweight alternative to state-of-the-art generative models. As an example, we compare train and inference times of EnScale to the best-performing and most recent ML-baseline, CorrDiff (Tab. 2). We find that training EnScale only takes about 1 day, whereas the diffusion model requires almost 10 times as long. The difference for inference is even larger, with EnScale being around 20 times faster. However, it should be noted that inference times can likely be improved for both models. For EnScale, the sampling can be done in a few minutes, but the current bottleneck is the inversion of the data-transformation.
EnScale is parameter-efficient due to its sparse layers (Sec. 2.4). Despite modeling effects specific to each location, the network only has 66 million parameters. The UNet in our CorrDiff implementation has around 80 million parameters (as in [Mardani2025ResidualDownscaling]) and is used in both the regression as well as the residual modeling step, yielding a total of around 160 million parameters.
Computational efficiency gains for EnScale are due to both the overall setup and the model architecture. The sparse local layers were designed for progressive resolution refinement in a stochastic manner, effective in combination with the energy score loss and our multi-step setup (Sec. 2.2), and are key to achieving the short training time. For example, training a fully convolutional architecture for the last steps of the super-resolution model took several days.
5.3 Exemplary samples
Figure 7: Exemplary samples from EnScale for all variables. The first column presents the unseen target RCM data by ALADIN63 driven by CNRM-CM5 on day 2035-05-06. Columns 2-4 show three corresponding random samples from EnScale. We chose a day with average performance, i.e., the EnScale-loss is roughly equal to the median score of all days in the interpolation test set.
Qualitatively, EnScale outputs realistic spatial fields that resemble the RCM data as illustrated in Fig. 7 for an exemplary GCM input. Examples for two other GCM inputs are shown in Fig. A.4 and A.5. For example, Fig. A.4 demonstrates a case where samples by EnScale are particularly similar to the RCM data (i.e., small EnScale-loss) .
The goal of EnScale is not to match the RCM data directly, but to generate plausible realizations given the GCM boundary conditions. We find that key features like topography and land-sea contrast are captured even though not provided explicitly as inputs. For temperature, samples align well with the RCM, with some differences in the Alps and in the eastern part of the domain, where variability is largest. For the other variables, the sampled fields broadly agree with the RCM, but exhibit substantial variability between samples for this particular GCM input. The samples differ both in location and intensity. Surface wind has a pronounced land-sea contrast, both in the RCM as well as in the samples, with higher magnitudes and variability over sea than over land. We further analyze and quantify sources for this variability in Sec. 6.
Per GCM-RCM pair, only one RCM field per GCM input is available. Thus, it is not possible to compare the variability in EnScale outputs with the raw data directly. Instead, one can follow the nearest-neighbor approach as in our analogue benchmark. Exemplary samples of this benchmark are shown in Fig. A.6. In this case, the variability stems partly from the small training dataset and imperfect near-neighbors (App. B.7). Nevertheless, the examples give a first impression of the variability that is present in the raw GCM-RCM data.
We present exemplary samples for the other machine learning methods (GAN and CorrDiff) in Fig. A.7 and A.8. We find similar conclusions as for EnScale: while the benchmark’s spatial fields resemble the RCM, they all exhibit considerable variability. GAN outputs tend to be blurry, whereas CorrDiff samples capture fine spatial details more realistically.
5.4 Evaluation of performance for individual variables
5.4.1 Overall performance with scoring rules
[…]
tas
Method MSE $\downarrow$ ES $\downarrow$ $_{\text{pred}} ES$$_{\text{var}}$ CRPS $\downarrow$ CRPS-mp $\downarrow$ RALSD $\downarrow$
NN-det 4.5 ± 0.08 245.5 ± 1.8 245.4 0.0 1.54 ± 0.01 1.51 ± 0.01 0.9 ± 0.01
EasyUQ 6.5 ± 0.08 182.3 ± 1.47 310.1 255.9 1.22 ± 0.01 2.25 ± 0.02 6.2 ± 0.02
Analogues 7.5 ± 0.12 241.2 ± 2.7 410.6 338.1 1.68 ± 0.01 1.63 ± 0.01 1.3 ± 0.01
GAN 9.2 ± 0.09 223.4 ± 1.44 376.1 305.4 1.48 ± 0.01 1.41 ± 0.01 2.4 ± 0.02
CorrDiff 6.5 ± 0.1 212.5 ± 2.21 396.6 369.0 1.09 ± 0.01 1.06 ± 0.01 1.0 ± 0.01
EnScale 4.6 ± 0.07 168.7 ± 1.59 356.4 375.2 1.20 ± 0.01 1.17 ± 0.01 1.4 ± 0.01
EnScale-t 5.2 ± 0.09 180.2 ± 1.84 367.1 373.9 1.26 ± 0.01 1.23 ± 0.01 1.4 ± 0.01
pr
method MSE $\downarrow$ ES $\downarrow$ $_{\text{pred}} ES$$_{\text{var}}$ CRPS $\downarrow$ CRPS-mp $\downarrow$ RALSD $\downarrow$
NN-det 30.9 ± 0.6 610.5 ± 5.72 610.0 0.0 2.33 ± 0.02 6.22 ± 0.06 10.1 ± 0.1
EasyUQ 49.5 ± 0.72 446.7 ± 4.48 817.2 740.6 1.94 ± 0.02 11.37 ± 0.07 9.6 ± 0.05
Analogues 39.6 ± 0.67 534.2 ± 5.96 892.3 715.9 2.35 ± 0.02 5.93 ± 0.05 5.2 ± 0.07
GAN 45.8 ± 0.9 460.5 ± 5.06 755.8 591.8 1.94 ± 0.02 5.16 ± 0.05 8.8 ± 0.07
CorrDiff 47.9 ± 0.78 414.7 ± 4.1 792.5 755.4 1.87 ± 0.02 4.68 ± 0.04 3.6 ± 0.05
EnScale 32.0 ± 0.61 434.5 ± 4.6 825.3 780.9 1.89 ± 0.02 4.85 ± 0.05 5.2 ± 0.06
EnScale-t 31.7 ± 0.62 433.7 ± 4.34 824.4 782.2 1.90 ± 0.02 4.86 ± 0.04 5.0 ± 0.06
sfcWind
Method MSE $\downarrow$ ES $\downarrow$ $_{\text{pred}} ES$$_{\text{var}}$ CRPS $\downarrow$ CRPS-mp $\downarrow$ RALSD $\downarrow$
NN-det 2.8 ± 0.03 204.8 ± 1.11 204.8 0.0 1.18 ± 0.01 1.50 ± 0.01 2.1 ± 0.02
EasyUQ 4.4 ± 0.03 147.3 ± 0.8 262.8 231.1 0.92 ± 0.01 2.25 ± 0.01 9.0 ± 0.02
Analogues 4.0 ± 0.04 183.0 ± 1.53 308.9 252.0 1.14 ± 0.01 1.41 ± 0.01 2.0 ± 0.02
GAN 4.9 ± 0.04 155.8 ± 0.88 275.4 239.3 0.97 ± 0.0 1.19 ± 0.01 3.7 ± 0.03
CorrDiff 4.0 ± 0.03 141.2 ± 0.86 276.5 270.8 0.85 ± 0.0 1.04 ± 0.01 1.5 ± 0.01
EnScale 2.8 ± 0.03 138.4 ± 0.92 273.9 271.1 0.89 ± 0.0 1.09 ± 0.01 1.8 ± 0.02
EnScale-t 2.9 ± 0.03 141.2 ± 0.9 274.3 266.3 0.90 ± 0.0 1.10 ± 0.01 1.9 ± 0.02
rsds
method MSE $\downarrow$ ES $\downarrow$ $_{\text{pred}} ES$$_{\text{var}}$ CRPS $\downarrow$ CRPS-mp $\downarrow$ RALSD $\downarrow$
NN-det 2034.3 ± 33.68 5040.2 ± 45.05 5039.9 0.0 31.06 ± 0.26 26.14 ± 0.23 3.8 ± 0.03
EasyUQ 3359.0 ± 39.28 3614.8 ± 30.22 6765.7 6296.6 24.28 ± 0.2 47.49 ± 0.3 10.3 ± 0.03
Analogues 2628.3 ± 42.57 4360.8 ± 47.03 7303.7 5897.6 28.69 ± 0.24 20.95 ± 0.18 1.8 ± 0.02
GAN 3052.0 ± 42.32 3640.3 ± 30.95 6380.9 5470.7 23.98 ± 0.19 17.77 ± 0.15 3.9 ± 0.03
CorrDiff 3638.9 ± 57.86 3181.2 ± 27.2 6473.6 6576.6 22.84 ± 0.2 15.98 ± 0.15 1.5 ± 0.01
EnScale 2060.0 ± 32.71 3478.2 ± 34.59 7142.6 7331.4 24.11 ± 0.19 21.56 ± 0.15 2.9 ± 0.02
EnScale-t 2084.1 ± 29.71 3495.7 ± 34.16 7158.7 7331.0 24.07 ± 0.21 21.53 ± 0.15 2.8 ± 0.02
a Metrics are first averaged across all days. Afterwards both metrics and standard errors are averaged across all GCM-RCM pairs to evaluate the overall performance on the entire dataset. Bold and underlined values highlight the best and second-best results, respectively.
Moving from visual evaluation to more quantitative results, Tab. 3 demonstrates the performance of EnScale in comparison to the various benchmarks explained in Sec. 4.1. To assess skill overall and in each location, we show MSE, energy scores (ES), and CRPS (see Sec. 4.2.1) separately for each variable. In general, we find that EnScale reaches good results compared to the benchmarks, often achieving the best or second-best results. EnScale-t shows similar results to EnScale, but usually performs slightly worse, since these metrics only evaluate the distribution for each day independently of the previous days. A dedicated evaluation of the temporal structure follows in Sec. 5.4.5. In the following, we analyze individual metrics in more detail.
The deterministic neural network excels regarding MSE for all variables. This is to be expected, as this approach explicitly minimizes the MSE during training. EnScale and EnScale-t perform only slightly worse. For all stochastic approaches (i.e., EasyUQ, analogues, GAN, CorrDiff, and EnScale), we approximate the conditional expectation $E^{\iota}[Y|X]$ as the average over nine samples corresponding to the particular GCM input $X$. With more samples per day, the estimate of the conditional mean improves, and lower scores are likely attainable, but the qualitative picture should persist.
When considering the energy score averaged over multiple GCM inputs $X$, CorrDiff or EnScale perform best, with the best method depending on the variable.
The energy score allows us to investigate the trade-off between prediction error ($\text{ES}_{\text{pred}}$) and variability ($\text{ES}_{\text{var}}$) quantitatively. If a method learned the correct RCM conditional distribution $Y|X$, these two scores should be equal ($\text{ES}_{\text{pred}}=\text{ES}_{\text{var}}$, see Sec. 2.5.3). For EnScale, we slightly underestimate variability for temperature and surface wind ($\text{ES}_{\text{pred}}>\text{ES}_{\text{var}}$), but overestimate variability for precipitation and solar radiation ($\text{ES}_{\text{pred}}<\text{ES}_{\text{var}}$). The latter could be a result of suboptimal convergence, whereas underestimation of variability can stem from mild overfitting. If $\text{ES}_{\text{pred}}=\text{ES}_{\text{var}}$ on the training data, $\text{ES}_{\text{pred}}$ usually increases on the test data, whereas $\text{ES}_{\text{var}}$ stays nearly constant, leading to $\text{ES}_{\text{pred}}>\text{ES}_{\text{var}}$ in test case.
We find that some other benchmarks (analogues, GAN) often severely underestimate the variability compared to the prediction error ($\text{ES}_{\text{pred}}>\text{ES}_{\text{var}}$). For applications, this would imply overconfidence, i.e., a lack of spread in the generated samples, which from a risk perspective is an undesirable property of a prediction system. For GANs, for example, the lack of variance is a well-known problem [Salimans2016ImprovedGANs]. For analogues, the relationship between $\text{ES}_{\text{pred}}$ and $\text{ES}_{\text{var}}$ depends on the number of near neighbors chosen, as alluded to in Sec. 4.1. In our setting, $k=5$ neighbors still underestimate $\text{ES}_{\text{var}}$.
Crucially, a lower prediction error ($\text{ES}_{\text{pred}}$) is not always desirable if it comes at the cost of reduced variability. For instance, NN-det has the lowest $\text{ES}_{\text{pred}}$, but their lack of variability leads to the poorest scores. Hence, if one is interested in the full conditional distribution, the goal is not only to minimize $\text{ES}_{\text{pred}}$ but to strike the right balance between prediction error and variability. We investigate this trade-off also visually, using the same example day that was presented above in Fig. 7. Fig. A.9 shows both the prediction error for one sample as well as the variability across several samples of EnScale. Both terms show similar magnitude and spatial structure, highlighting that EnScale balances prediction error and variability terms. For comparison, Fig. A.10 shows the same for the analogues, with qualitatively similar conclusions.
Furthermore, we assess distributions in each location separately with the CRPS. CorrDiff reaches best scores for all variables, closely followed by EnScale. Comparison with the EasyUQ benchmark serves as another sanity check: EasyUQ is expected to perform well as it explicitly aims to best capture the pointwise CRPS values. EnScale reaches slightly better scores than the EasyUQ benchmark, showing that we successfully learned the distributions in each location despite the high-dimensional target.
[…]
Scores can differ considerably between different pairs. For example, energy scores for temperature vary between 131 and 193, and those for precipitation between 359 and 562. This indicates that various RCMs disagree on the underlying variability, or possibly are not equally easy to emulate with the ML model. All above tables only consider the interpolation test set, i.e., the held-out years 2030-2039. We provide scores on the extrapolation test set (2090-2099) (see Sec. A.4). Scores are very similar and the conclusions are consistent.
5.4.2 Spatial structure
One way to assess the spatial structure is by pooling pixels spatially and computing the CPRS on these pooled results. Tab. 3 shows the results where we applied max pooling across patches of size $10\times 10$ before calculating CRPS. These findings remain consistent for other kernel sizes for pooling. The benefit of our spatially multivariate approach becomes apparent in these results. The EasyUQ benchmark, in contrast to good pointwise CRPS scores, fails to capture correct spatial structures and leads to large values for this metric. CorrDiff, EnScale and EnScale-t again show good performance and are superior to the other benchmarks.
In addition, we compute power spectral densities (PSDs) for all days and generated samples in the test set and report radially-averaged log spectral distance (RALSD, see Sec. 4.2.2). CorrDiff reaches best results, except for temperature, where NN-det has lowest RALSD. EnScale or analogues are second-best. Performance with RALSD depends on both the spatial quality as well as the similarity of samples and truth. Alternatively, we average power spectra across multiple days before calculating log spectral distances (RALSD-avg, see Sec. 4.2.2), results are shown in Tab. A.3. Here, the analogues reach lowest errors as expected because they reproduce correct spatial fields.
Plots for the PSDs in Fig. A.14 confirm that both CorrDiff and EnScale closely matches power spectra of the RCM. We find only a slight overestimation of the power for precipitation and solar radiation on small scales. Other results are as expected: EasyUQ’s have too much small-scale spatial structure, overestimating the power at small wavelengths and underestimating it at larger wavelengths. NN-det’s outputs are too smooth, generally underestimating the PSDs across wavelengths, most strikingly for precipitation and wind.
5.4.3 Calibration
Figure 8: Rank histograms for evaluating calibration. For each day, we calculate the spatial mean and the spatial maximum for the RCM data and for 9 samples of each method, and construct rank histograms for these spatial statistics (first and second rows show the spatial mean, third and fourth row the spatial max.). The flat dashed line represents the uniform distribution, indicating perfect calibration. Miscalibration (MCB) scores in the legend quantify deviations from the uniform distribution at each individual grid point, averaged across all locations.
We assess the calibration of the samples with rank histograms (see Sec. 4.2.3). First, we analyze calibration of spatially-aggregated quantities, the spatial mean and maximum, in Fig. 8. We find that most approaches fail to achieve good calibration for most variables, with varying degrees of violation and biases. In general, calibration for the spatial mean is often better than that for the spatial maximum.
Considering the spatial mean, we find that EnScale only shows mild violations from calibration. The samples tend to underestimate the RCM’s values for precipitation and surface wind (higher counts for larger ranks), and are slightly overdispersed for solar radiation (dome-shape of the rank histogram). However, the benchmarks do not perform better: CorrDiff shows large deviations from calibration. We find that analogues are usually underdispersed, which is in line with the scores in Tab. 3 ($\text{ES}_{\text{var}}<\text{ES}_{\text{pred}}$), and the GAN also shows errors in dispersion or biases.
The spatial maximum is more difficult to capture for some variables. Most approaches are not calibrated. For EnScale, we observe underestimation of the spatial maximum for precipitation and overestimation of that for temperature, wind and solar radiation.
We compare the average miscalibration metric (MCB) across locations (see Sec. 4.2.3), see scores in the legend of Fig. 8. EnScale and EnScale-t show errors in calibration with MCB values between 0.09 and 0.17, outperforming CorrDiff. Also analogues and GAN reach higher MCB scores than EnScale (not shown). For reference, MCB scores for the spatial mean vary between 0.08 (temperature) and 0.2 (precipitation). Locationwise calibration is generally best for temperature and wind and worst for solar radiation and precipitation.
Tracing down the root of the calibration issues in EnScale is challenging due to the multi-step approach. We find that the coarse model is close to calibration for the spatial mean, but the issues for the spatial maximum are already evident in this step. The several steps of the super-resolution model show deviations from calibration for both the spatial mean and maximum.
These results show room for improvement for all ML approaches and point to limitations of the energy score as a loss function. In principle, a method that reaches the minimal energy score also has calibrated samples. However, in practice, the energy score is not equally sensitive to all aspects of misspecification [Pinson2013DiscriminationScore], and in some cases, a slightly sharper yet miscalibrated forecast may achieve a better energy score than a well-calibrated but less sharp one. This finding is in line with other works in weather post-processing [Wilks2018EnforcingPostprocessing, Wessel2025EnforcingModels]. It suggests that additional measures may be needed to explicitly assess and improve calibration in generative approaches.
5.4.4 Extremes
Table 4: Extreme evaluation regarding marginal quantiles as well as miscalibrationa
Method tas pr
AE $Q^{0.05}_{\text{Wi}}\downarrow$ AE $Q^{0.95}_{\text{Su}}\downarrow$ MCB $Q^{0.1}$ $\downarrow$ MCB $Q^{0.9}$ $\downarrow$ AE $Q^{0.05}_{\text{Wi}}\downarrow$ AE $Q^{0.95}_{\text{Su}}\downarrow$ MCB $Q^{0.1}$ $\downarrow$ MCB $Q^{0.9}$ $\downarrow$
NN-det 1.021 0.948 nan nan 0.007 6.960 nan nan
EasyUQ 0.539 0.479 0.060 0.064 0.000 2.459 0.028 0.039
Analogues 0.847 0.485 0.103 0.054 0.000 1.449 0.068 0.085
GAN 1.606 1.545 0.097 0.063 0.012 4.395 0.274 0.038
CorrDiff 1.069 1.086 0.056 0.050 0.044 5.327 0.088 0.047
EnScale 0.488 0.478 0.021 0.021 0.000 2.813 0.045 0.018
EnScale-t 0.673 0.817 0.025 0.018 0.000 2.869 0.047 0.019
Method sfcWind rsds
AE $Q^{0.05}_{\text{Wi}}\downarrow$ AE $Q^{0.95}_{\text{Su}}\downarrow$ MCB $Q^{0.1}$ $\downarrow$ MCB $Q^{0.9}$ $\downarrow$ AE $Q^{0.05}_{\text{Wi}}\downarrow$ AE $Q^{0.95}_{\text{Su}}\downarrow$ MCB $Q^{0.1}$ $\downarrow$ MCB $Q^{0.9}$ $\downarrow$
NN-det 0.457 0.929 nan nan 6.744 15.976 nan nan
EasyUQ 0.108 0.211 0.041 0.049 1.333 3.178 0.023 0.038
Analogues 0.102 0.254 0.076 0.095 1.110 3.730 0.064 0.119
GAN 0.279 0.775 0.064 0.068 7.620 13.761 0.091 0.038
CorrDiff 0.282 0.491 0.059 0.043 3.219 9.879 0.051 0.097
EnScale 0.118 0.278 0.013 0.015 2.623 5.488 0.017 0.024
EnScale-t 0.124 0.266 0.013 0.025 2.377 5.597 0.015 0.023
a For each method, we compute extreme quantiles marginally across all days for two seasons (winter / summer). We report absolute errors (AE) compared to RCM data (“AE Q” columns). The miscalibration metric for quantiles (“MCB Q” columns) evaluates reliability of quantiles in the conditional distribution of the samples given GCM input.
We focus on the capability of EnScale and the benchmark to capture extreme events. Firstly, we assess locationwise quantiles in Tab. 4. EnScale performs well compared to the benchmarks, but best scores are often reached by the analogues and EasyUQ. We note, however, that these estimates are conservative – per season, there are approximately $1000$ days in the test set, leading to estimation errors for extreme quantiles. Notably, the GAN and CorrDiff struggle to reproduce extreme quantiles accurately. In particular, they miss to predict the zero precipitation events correctly. The true 5 % -quantile in winter is zero, but the ML approaches do not capture this. EnScale successfully predicts zero rain events. This is an advantage of our data transformation (Sec. 3.2).
Figure 9: Extreme quantiles for the summer season (June, July, August) compared to ALADIN63 driven by CNRM-CM5. The first column shows the 95 % - quantile of the RCM calculated separately in each location. The remaining columns show the error for different methods, calculated as the signed difference of quantiles in the samples minus the RCM’s quantiles. Additionally, the last column provides an assessment of “model uncertainty” where the difference of ALADIN63 to CCLM4-8-17 (both driven by CNRM-CM5) is shown. Similar results are obtained for pairs other than ALADIN63 driven by CNRM-CM5.
A more detailed analysis of the 95 % - quantile in summer is shown in Fig. 9. This illustration reveals the direction of error, which varies across locations and variables. For example, EnScale overestimates extreme temperatures but underestimates extreme solar radiation in the majority of locations. For precipitation, absolute errors are highest in the Alpine region, and for surface wind, errors are largest over the Sea. Again, EnScale performs well in comparison to the benchmarks, with errors of similar magnitude as for the analogues and often lower than for GAN and CorrDiff.
In order to provide an additional perspective on these results, we compare to uncertainties across different RCMs. As an example, we show differences between the quantiles of two RCMs, both driven by the same GCM. We find that the mismatches between RCMs are of similar magnitude to the errors of the ML models. Importantly, in most cases, disagreements between the RCMs exceed the errors of EnScale. This suggests that, despite some limitations, the emulators typically produce results that fall within the range spanned by multiple RCMs.
Next, we consider calibration of low / high quantiles of the predicted conditional distribution (see “MCB Q” columns in Tab. 4 and Sec. 4.2.4 for details). Here, EnScale and EnScale-t usually reach the lowest errors, with the exact ordering of these three methods depending on the variable and quantile considered. This again exemplifies that EnScale is sometimes inferior to the analogues with respect to metrics evaluating the marginal distributions $p_{Y}$. For the conditional $p_{Y|X}$, however, EnScale is superior.
5.4.5 Temporal structure
Table 5: Error in lag-1 autocorrelations (ACF) for each variablea
Method tas pr sfcWind rsds
NN-det 0.01 0.17 0.08 0.11
EasyUQ -0.03 -0.07 -0.15 -0.03
Analogues -0.10 -0.12 -0.28 -0.04
GAN -0.06 -0.1 -0.21 0.02
CorrDiff 0.01 0.23 0.10 0.08
EnScale -0.08 -0.09 -0.22 -0.06
EnScale-t -0.01 -0.01 -0.04 -0.01
aWe calculate ACFs in each location for both the samples and the RCM data.We average the signed difference of samples’ ACF minus target ACF over all locations.
We investigate the temporal structure of the generated samples, that is, how accurately the methods capture correlation between consecutive time steps. To this end, Tab. 5 shows differences in lag-1 (i.e., a single time step) autocorrelations (ACFs) between the samples and the RCM data.
We find that all methods except NN-det and CorrDiff have negative differences, i.e., they underestimate the local autocorrelation. CorrDiff accurately reproduces autocorrelations for temperature, but overestimates them for all other variables (see App. A.2 for more details). EnScale-t is able to reduce the error substantially and has only a small tendency to underestimate the autocorrelation for sfcWind. To ensure that the error reduction does not stem from canceling errors due to opposite signs, we compare absolute errors in Tab. A.2. We find that EnScale-t is superior to the others and achieves the lowest deviation from the true ACFs across all variables.
Capturing autocorrelations is particularly relevant when analyzing statistics aggregated over multiple days, such as in 5-day heatwaves. Underestimating autocorrelation can lead to an underrepresentation of variability in multi-day averages, therefore potentially underestimating risk. These findings point to limitations of existing ML-based downscaling approaches that fail to include temporal correlation on a local scale. EnScale-t has small room for improvement, especially for wind over the Alps (not shown). This is likely due to the approximation of modeling temporal consistency only on the level of the coarsened RCM (see Sec. 2.3). Nevertheless, even with this approximation, we can see a clear benefit compared to the other models.
Figure 10: As Fig. 6, omitting CorrDiff, but presenting all four target variables instead.
We show an exemplary time series in Fig. 10 for the spatial mean over one year. We compare EnScale, EnScale-t, and the underlying RCM. Most importantly, we find that generated time series by both, EnScale and EnScale-t, closely match the ground truth over the entire year and capture the seasonal cycle. On most days, the truth lies within the range of the 9 samples.
The illustration highlights the stability of the time series of the autoregressive rollout for our temporally-consistent coarse model. Both the individual time series as well as the uncertainty range remain stable over time. We also tested longer rollout of our generated time series and found that they are stable even for several years (not shown). Both the time series as well as the loss do not increase over time. The stability of rollout is an open discussion in the case of ML-based weather forecasting [Price2025ProbabilisticLearning, Karlbauer2024AdvancingMesh]. In our case, the additional input from the GCM likely helps to stabilize the rollout and prevent potential drifts.
The insets in the plots highlight qualitative differences between generated time series by EnScale-t and EnScale. The time series for the EnScale-t tend to be smoother because they take into account the temporal structure, especially for temperature. The time series by EnScale is not necessarily always consistent over time as each day is generated independently of the previous. The uncertainty bands are similar for both approaches. In times where temporal autocorrelation is high (especially visible for temperature), the uncertainty range for the EnScale-t is larger because including temporal structure allows us to sample a larger variety of possible trajectories.
5.5 Variable dependencies
Table 6: Multi-variable evaluationa
Method Correlations Miscalibration
pr-sfcWind sfcWind-rsds tas-pr pr-sfcWind sfcWind-rsds tas-pr
NN-det 0.125 0.062 0.060 nan nan nan
EasyUQ 0.166 0.217 0.055 0.563 0.790 0.389
Analogues 0.026 0.029 0.025 0.355 0.331 0.325
GAN 0.063 0.056 0.042 0.098 0.092 0.123
CorrDiff 0.126 0.094 0.117 0.195 0.192 0.213
EnScale 0.044 0.029 0.031 0.055 0.054 0.067
EnScale-t 0.042 0.031 0.033 0.056 0.054 0.067
aWe calculate pairwise correlations between two variables in both the samples and the RCM data, and report absolute errors averaged over 1000 randomly chosen locations (first three columns). The last three columns show miscalibration of the pairwise differences of variables. Bold and underlined values highlight the best and second-best results, respectively.
So far, our evaluation considered each target variable separately. Next, we evaluate the capability of our model and the benchmarks to capture the full distribution jointly for the four targets in Tab. 6.
Firstly, we consider pairwise correlations between variables for three selected variable pairs. For all pairs, the analogue benchmark reaches best or second-best scores. This is because the analogues only reproduce existing RCM data. By construction, the dependencies between variables are correct for the existing data. Results by EnScale are similar. In contrast, EasyUQ which models each variable independently of others, shows much higher errors. This is another example for the benefits of our multivariate approach.
Figure 11: Correlations between pairs of variables. The first column shows the RCM’s pairwise correlation between the variables in each location. The other columns show the error of EnScale and other methods, i.e., the signed difference of correlations in the samples minus the RCM. This figure is based on the results for CNRM-CM5 and ALADIN63 only, but other pairs reveal a similar pattern. The last column “model uncertainty” refers to the difference in correlations of two example RCMs, ALADIN63 and CCLM4-8-17, both driven by CNRM-CM5.
Fig. 11 presents the pairwise correlations in more detail. The patterns for the RCM show local details, especially in mountain regions. Errors for EnScale and EnScale-t are small, comparing favorably to the benchmarks. Sign and magnitude of errors vary across space. As in Sec. 5.4.4, we again put ML model errors into context to RCM disagreements. We present differences between the inter-variable correlations of two RCMs. Our findings are similar to what was observed previously: RCM differences are about the same magnitude as ML model errors, with EnScale often reaching smaller errors than the RCM mismatches.
The next metrics evaluates dependencies of the variables conditional on GCM data $X$: we consider calibration of the pairwise difference of variables (last three columns of Tab. 6, see Sec. 4.2.6 for details). Now, performance of the analogues is worse. Additional evaluation shows that they are again underdispersed (not shown). Instead, EnScale or EnScale-t achieve the best results. This finding again highlights the difference when analyzing marginals vs. conditional distributions, as in Sec. 5.4.4.
We note that the above metrics only evaluate pairwise distributions between variables. Assessing the full joint distribution is not considered. In addition, we don’t explicitly evaluate physical consistency of variable dependencies. However, as the RCM’s output is physically consistent, similarity of the ML’s samples to the RCM implicitly evaluates this. Future work could focus more on physics-based evaluation.
6 Variability in EnScale’s samples
The conditional variance for the full GCM-RCM path is large, as can be seen from diversity in the emulator’s sampled fields (Fig. 7). This variability arises from internal variability in the RCM given the GCM’s boundary conditions (see Sec. 5.3). As each RCM downscales each GCM input only once, we cannot quantify this variability in the RCM data directly. Instead, we use EnScale to analyze this spread. Firstly, we split the variance of the full path from GCM to RCM into variance on coarse scales (path from GCM to coarsened RCM) and on local scales (path from coarsened RCM to RCM). Secondly, we compare internal variability with structural differences between RCMs.
6.1 Coarse-scale vs. local-scale variability
Table 7: Variance decomposition coarse scale vs high-resolution scalea
Variable Coarse-scale variance Local-scale variance Ratio (coarse/local)
tas 3.97 0.24 16.76
pr 17.15 8.91 1.96
sfcWind 1.86 0.31 6.05
rsds 1387.64 287.63 4.81
aWe split the full variance in EnScale’s samples into a contribution stemming from the variance of the coarse model (first column), and one of the super-resolution model (second column). The decomposition is formalized in Eq. (4) in the methods. The last column shows the ratio of coarse / local, indicating the level of variance that is present on the coarse compared to the high-resolution scale. Variances are given in the square of the native units ($\mathrm{K}^{2}]$ for temperature, $\mathrm{(mm/day)}^{2}]$ for precipitation, $\mathrm{(m/s)}^{2}]$ for surface wind, $\mathrm{(W/m^{2})}^{2}]$ for solar radiation), and the ratio is dimensionless.
EnScale naturally allows us to separate sources of variability on different scales, as it splits the full GCM-RCM map into two steps. First, the coarse model maps GCM predictors to average-pooled RCM data and second, the super-resolution model learns the distribution of the high-resolution target given the average-pooled predictors (Fig. 2). This means we can split the full variability into contributions from the two paths: (i) coarse-scale variability, quantifying uncertainties in the pooled RCM data given GCM inputs, and (ii) local-scale variability, capturing variance in high-resolution RCM fields when conditioned on their coarse averages (see Sec. 4.3.1 for the formal variance decomposition).
For a first qualitative insight, we first compare EnScale’s full GCM-RCM map (as in the examples in Fig. 7) with examples from the pure super-resolution model. Using average-pooled versions of the target RCM data as inputs, EnScale’s samples agree very closely with the target (see Fig. A.11). The different samples show only small discrepancies on a local scale and variability is greatly reduced in this case compared to the full GCM-RCM model. This highlights that the full RCM emulation is a very different task from a pure super-resolution setup. In addition, it indicates that variability between GCM and RCM on coarse scales is a main source of the mismatches that we observed in samples by EnScale (see Fig. 7) and benchmarks (see Fig. A.6A.7).
The quantitative result of the decomposition is presented in Tab. 7, for each variable separately, averaged over locations. For all variables, the ratio of coarse-scale to local-scale variance greatly exceeds 1, indicating that most variability stems from the transfer of GCMs to average-pooled RCMs. In other words, the GCM boundary conditions contain comparatively little information about the coarsened RCM data, whereas the coarsened RCM already well determines the high-resolution output.
The ratio is between 2.0 and 16.8 for the different variables. It is largest for temperature and wind and smallest for precipitation. For wind, the high ratio is primarily due to large-scale variability over sea areas, where local variance is small. In contrast, precipitation exhibits greater local-scale variability, particularly in the Alpine regions. Since these small-scale effects cannot be fully captured by the average-pooled data $Z$, the contribution of local-scale variance remains relatively large, leading to a lower overall ratio.
Individual variances can be interpreted. We consider temperature, which is approximately normally distributed. Calculating $\pm 2\sqrt{\operatorname{Var}}$ gives the range around the mean in which approx. 95 % of data points lie. For example, for the local-scale variance $\pm 2\sqrt{\operatorname{Var}}=\pm 0.98$. That means, after fixing the average-pooled RCM data, local temperatures fluctuate by around $\pm$ $1\text{\,}\mathrm{K}$ in each grid point. Conditioning on the GCM input, however, temperatures on coarse scales can still vary by about $\pm$ $4.0\text{\,}\mathrm{K}$.
Local-scale variance in the RCM is to be expected, as the average values over larger areas can never determine the high-resolution output uniquely. Nevertheless, these findings indicate that variability on coarse scales is even more prominent. The GCM input gives only small constraints on the average-pooled RCM data. We suspect that this is because the GCM only drives the RCM at the boundaries, leaving a wide range of plausible climate conditions inside the RCM’s domain.
6.2 Internal variability vs. model differences
Table 8: Variance decomposition for different RCMs with the same driving GCMa
Variable Variance of RCM Model difference Ratio (Variance / model diff.)
tas 4.42 0.42 10.50
pr 27.67 1.25 22.17
sfcWind 2.08 0.14 14.97
rsds 1871.04 294.08 6.36
aUsing EnScale, we compare variance within samples for individual RCMs to the structural difference between RCMs, following the decomposition derived in Eq. (5) in the methods. Variances are expressed conditional on GCM inputs $X$, and averaged over all inputs in the test set. We use EnScale’s samples for ALADIN63 as a baseline run and compare to samples for CCLM4-8-17 (both driven by CNRM-CM5). The last column shows the ratio the variance of RCMs vs. the model differences. The units for the variance and model differences are again squares of the native units, as in Tab. 7, and the ratio is dimensionless.
We use EnScale to analyze internal variability relative to model disagreements. Recall that EnScale emulates the distribution of each RCM specifically ( see Sec. 2.2). We compare variability within EnScale’s samples for individual RCMs to differences between samples for two RCMs, as described in Sec. 4.3.2, in particular Eq. (5). Tab. 8 presents the results.
We find that variances within samples for an individual RCM largely dominate over differences between samples for two distinct RCMs (i.e., ratio larger than 1). The ratio is largest for wind, where the variance is high for EnScale’s samples for each RCM alone already. We show results for samples for two RCMs only, both with the same GCM input. Overall findings are similar also for other pairs.
The variances here can be interpreted as well, similar to the discussion in Sec. 6.1. For example, for temperature, the variance of one RCM is $3.9\text{\,}\mathrm{K}^{2}$. That means, for a fixed GCM input, temperatures simulated by the RCM in each location can fluctuate by $\pm$4.2\text{\,}\mathrm{K}$\$ (based on an approximation with a normal distribution, and calculated as above as $2\sqrt{\operatorname{Var}}$). However, the disagreements between RCMs are much smaller than that. In each location, the means of the RCMs differ by about $1.3\text{\,}\mathrm{K}$, when conditioning on a GCM input (calculated as $2\sqrt{\text{model diff.}}$, compare with the formula (5) in Sec. 4.3.2).
This finding points to the fact that internal variability seems to dominate over structural model differences for regional climate model ensembles. It is similar to [Rampal2025DownscalingExtremes], who use an RCM-emulator to present an uncertainty quantification into internal variability, scenario uncertainty and model uncertainty and also find a prominent role of internal variability. In addition, our variance decomposition underlines the benefits of jointly training one ML model for downscaling based on multiple GCM-RCM combinations. Even though inter-model differences can be rather large, intrinsic conditional variance of $Y|X$ dominates. Hence, there is sufficient similarity between GCM-RCM pairs such that the ML training can benefit from a joint training.
7 Discussion
In this study, we present EnScale, a method for emulating RCMs. We compare EnScale to various benchmarks, both physically-motivated (analogues) and existing state-of-the-art machine learning methods (GAN and a diffusion model, CorrDiff). We find that EnScale achieves great performance in terms of a wide range of metrics, both when evaluated per grid point and per variable separately, as well as for spatial structure and dependencies between variables. To our knowledge, EnScale-t is the first downscaling approach that models temporal dependencies in the high-resolution output, improving the temporal consistency of the output substantially compared to all other benchmarks.
Compared to physics-based climate models as well as state-of-the-art machine learning models, our emulator is computationally very efficient. Training EnScale completes in approximately one day. Afterwards, generating a downscaled ensemble with ten members for 100 years of GCM data is completed in under two hours. CorrDiff instead needs about 10 days for training and 36 hours for inference. Running the RCM for the same period would typically take weeks and still yield only one ensemble member. The sparse architecture in EnScale is efficient and scalable also for larger domains and more climate variables.
Our work established a comprehensive evaluation framework of generative downscaling approaches for calibration, spatial and temporal structure, extremes, and multivariate dependencies. Notably, evaluating such methods is not a straightforward task. The output comprises many locations, variables and time steps, and capturing all these aspects is challenging [Maraun2015VALUEStudies]. Model evaluation and statements about skill are necessarily specific for each question and application. So far, our evaluation methods are primarily statistically motivated and inspired by the field of forecast evaluation. There remain open questions, in particular for evaluating the dependencies for many variables at once and for the temporal structure. Evaluating emulators in application-relevant case studies is an interesting avenue of further research [Isphording2024ASimulations].
We found large conditional variability in RCM fields given GCM inputs, arising from internal variability within RCMs. This highlights the value of stochastic emulators: by generating an ensemble of many RCM trajectories for one given driving GCM run, they can capture this internal variability and characterize its spread. This can enable better uncertainty quantification in high-resolution predictions, analogous to the benefits of large ensembles from global climate models [Deser2020InsightsProspects, Rampal2025DownscalingExtremes].
EnScale can readily be applied to data from the three GCMs that were used during training. It can also help to fill the GCM-RCM data matrix, provided that individual GCMs and RCMs were present in the training data. For example, it is possible to sample data from MIROC5-ALADIN63, even though this pair was not present in the training data. We also plan to provide downscaled CMIP6 data after re-training EnScale on more GCMs in CMIP5.
Our downscaling via coarse correction is similar to some existing works: [PriceStephanRasp2022IncreasingModels] also split downscaling of precipitation with a GAN into two steps: a coarse correction step and a pure super-resolution step. The target is radar measurements. Their setup is simpler than ours because they only use data from one numerical model, but still they find that isolating the correction step in a low-dimensional task improves the results. [Wan2024StatisticalModels] downscale data from a single GCM stochastically, considering historical GCM data and using ERA5 reanalysis as a target. Hence, their first step corrects large-scale mismatches between GCMs and ERA5. In a similar fashion, CorrDiff’s mean-residual split (Sec. 4.1.5) corrects some mismatches between input and target data in a simpler task (the mean prediction, first step), before learning the correct variability (residual prediction, second step).
We choose to learn the full path from GCM to RCM, which differs from RCM-emulation in [Doury2024OnPrecipitation, Doury2022]. Previous work argues that emulators which learn a map from GCM to RCM data might pick up patterns specific to each GCM-RCM pair, and are thus less transferable to other GCMs [Hernanz2023OnIssues, Rampal2024EnhancingLearning, Bano-Medina2023TransferabilityApplications]. However, our setup is different because we (i) train with multiple GCM-RCM pairs, possibly enabling the model to learn shared properties across different climate models, and (ii) train a generative model rather than a deterministic one. Future work could also include extrapolation to unseen GCMs.
EnScale leverages both past and future data, enabling it to learn future relationships between low-resolution predictors and high-resolution targets directly from physics-based models. In contrast, model training in perfect prognosis downscaling relies on historical data and aims to extrapolate these learned relationships to future conditions. RCM emulation like in EnScale can readily be combined with a state-of-the art bias-correction method to obtain “future observations”. A promising direction for further development is to explore more sophisticated bias correction algorithms that preserve multivariate structure across time, space, and several variables or integrate observational data also in model training [Hess2025FastLearning, Schmidt2025ASimulation].
EnScale is computationally efficient and can be re-trained also on other datasets from other GCMs and RCMs or for other regions. Also, our emulator could be used to downscale to a higher resolution, for example approximating convection-permitting models as a target. These benefits establish EnScale as a powerful tool for data-driven future climate projections on a local scale and open new opportunities for climate impact studies, where reliable high-resolution projections are essential.
Acknowledgments
Maybritt Schillinger, Maxim Samarin, Nicolai Meinshausen, and Reto Knutti are part of SPEED2ZERO, a Joint Initiative co-financed by the ETH Board. We thank Ruth Lorenz and Urs Beyerle for their help with the data-preprocessing and Sebastian Sippel for valuable inputs during the early stages of this project. Further, we benefitted from many helpful and inspiring discussions with numerous people, among them were Antoine Doury, Jose González-Abad, Shahine Bouabid, Sam Allen, Anna Merrifield, Dominik Schumacher, Daniele Nerini, Niklas Boers and his group. Last but not least, we thank Nina Effenberger and Anya Fries for their constructive feedback on the draft which greatly improved the readability of this work.
Code Availability
Appendix A Additional results
A.1 Distribution shifts between different RCMs
Figure A.1: As Fig. 5, but only comparing to EnScale-no-label and EnScale-t-no-label which exclude the one-hot-encoded index of the GCM and RCM also in the super-resolution model.
We trained two versions of the super-resolution model from coarsened RCM data to the high-resolution target. Once, we included a one-hot-encoded index for the RCM as an additional predictor (as in EnScale, EnScale-t), and once we excluded the index (version EnScale-no-label, EnScale-t-no-label). The index allows the ML model to adjust the output accordingly, dependent on the distribution of this particular RCM. Fig. A.1 summarizes the performance of EnScale-no-label, EnScale-t-no-label compared to EnScale. We find that adding the label only has a very small effect on the results. EnScale often yields small improvements compared to EnScale-no-label, in particular for extremes. This indicates that the step from coarsened RCM data to RCM data is indeed very similar between the different RCMs. For comparison, for the step from GCM to coarsened RCM, including the indices of the RCM improves the energy scores loss by about 8 %.
A.2 Comparison of CorrDiff and CorrDiff-s
CorrDiff (following [Mardani2025ResidualDownscaling] and the implementation in NVIDIA’s PhysicsNeMo library [^3] Code: https://github.com/NVIDIA/physicsnemo/tree/main/examples/weather/corrdiff) generates each ensemble member by drawing a single random noise realization that is shared across all time points. This construction implicitly imposes temporal dependence in the generated time series, but variability across realizations is limited. For this reason, we also tested an adapted version which we call CorrDiff-s. In this variant, we randomize the sampling scheme: For each day, noise is drawn independently of previous days and of other ensemble members. This version represents the variance of samples more faithfully. However, it removes temporal autocorrelation, leading to negative errors in autocorrelations, similar to the other benchmarks (not shown).
CorrDiff-s outperforms CorrDiff in almost all categories (Fig. A.2). As one example, we show that rank histograms by CorrDiff-s are much closer to calibrated (Fig. 8).
We suspect that these discrepancies are a result of large conditional variance in the distribution of RCM fields given GCM inputs. In this case, the original CorrDiff sampling scheme is suboptimal, whereas an independent sampling strategy better captures the spread of possible high-resolution fields.
Figure A.2: As Fig. 5, but comparing CorrDiff, CorrDiff-s, EnScale and EnScale-t.
Figure A.3: As Fig. 8, but comparing CorrDiff, CorrDiff-s, EnScale and EnScale-t.
A.3 Further exemplary samples
We present further examples for EnScale in Fig. A.4 and Fig. A.5, for the benchmarks in Fig. A.6 – Fig. A.8. In addition, we present the balance of prediction error and variability in Fig. A.9 and Fig. A.10. Finally, Fig. A.11 presents results for EnScale’s super-resolution model.
Figure A.4: Examples for EnScale: Like Fig. 7, but showing a day with particularly good scores (about 10-th percentile).
Figure A.5: Examples for EnScale: Like Fig. 7, but showing days with particularly bad scores (about 90-th percentile).
Figure A.6: Examples for the analogues: Like Fig. 7, but showing examples for the analogue benchmark.
Figure A.7: Examples for the GAN: Like Fig. 7, but showing examples for the GAN benchmark.
Figure A.8: Examples for CorrDiff: Like Fig. 7, but showing examples for the CorrDiff benchmark.
Figure A.9: Balance of prediction error vs. variability for EnScale: The first column shows the prediction error for one of the samples from EnScale, i.e. the difference of the generated sample to the RCM truth. Columns 2-4 show the variability, i.e. difference between three pairs of EnScale-samples.
Figure A.10: Like Fig. A.10, but for the analogue benchmark.
Figure A.11: Examples for pure super-resolution map: Like Fig. 7, but instead of presenting the full map from GCM data $X$ to RCM data $Y$, we show the results of only the super-resolution map from coarsened RCM data $Z$ to RCM $Y$.
A.4 Results on extrapolation test period
We present a selection of results also on the extrapolation test period (2090-2099). Fig. A.12 summarizes the evaluation and Tab. A.1 (as Tab. 3 in the main text) presents performance w.r.t. MSE, energy score, CRPS per location and after max-pooling as well as RALSD. The ordering of the methods remains consistent with the interpolation period, and absolute scores are very similar. To directly compare scores on interpolation and extrapolation test set, we compute the ratios between the two test periods (Fig. A.13). We find that most ratios are close to one, indicating at most small differences in performance between the two periods. Also, different methods seem to extrapolate similarly well. This is expected here since this is only a mild extrapolation case: the training data extend until 2089, and the test period (2090–2099) differs only slightly in climate conditions, with long-term climate change signals being small compared to daily variability.
Figure A.12: As Fig. 5, but showing results for the extrapolation test period.
Figure A.13: As Fig. 5, but comparing performance on extrapolation vs. interpolation test set: For each method and each metric, we calculate the ratio of the metric on the extrapolation test set vs. that on the interpolation test set.
Table A.1: As Tab. 3, but for the extrapolation test set.
tas
method MSE $\downarrow$ ES $\downarrow$ $_{\text{pred}} ES$$_{\text{var}}$ CRPS $\downarrow$ CRPS-mp $\downarrow$ RALSD $\downarrow$
NN-det-extra 4.4 ± 0.08 240.1 ± 1.9 240.0 0.0 1.51 ± 0.01 1.51 ± 0.01 0.8 ± 0.01
EasyUQ-extra 6.3 ± 0.08 179.7 ± 1.5 305.4 251.3 1.20 ± 0.01 2.14 ± 0.02 6.1 ± 0.01
Analogues-extra 8.0 ± 0.13 248.8 ± 2.7 415.5 332.8 1.74 ± 0.01 1.73 ± 0.01 1.2 ± 0.01
GAN-extra 9.3 ± 0.1 223.4 ± 1.42 377.4 308.1 1.48 ± 0.01 1.45 ± 0.01 2.2 ± 0.01
CorrDiff-extra 6.8 ± 0.11 219.5 ± 2.4 393.7 348.3 1.09 ± 0.01 1.08 ± 0.01 0.9 ± 0.01
EnScale-extra 4.4 ± 0.07 166.8 ± 1.57 348.5 363.5 1.18 ± 0.01 1.16 ± 0.01 1.3 ± 0.01
EnScale-t-extra 5.3 ± 0.09 178.5 ± 1.82 362.1 367.2 1.27 ± 0.01 1.26 ± 0.01 1.3 ± 0.01
pr
method MSE $\downarrow$ ES $\downarrow$ $_{\text{pred}} ES$$_{\text{var}}$ CRPS $\downarrow$ CRPS-mp $\downarrow$ RALSD $\downarrow$
NN-det-extra 35.1 ± 0.84 631.7 ± 6.64 632.3 0.0 2.38 ± 0.02 6.41 ± 0.07 10.5 ± 0.13
EasyUQ-extra 54.3 ± 0.92 470.0 ± 4.89 838.7 737.2 1.98 ± 0.02 11.39 ± 0.08 9.6 ± 0.05
Analogues-extra 45.7 ± 0.91 567.0 ± 6.95 936.3 738.4 2.43 ± 0.02 6.29 ± 0.06 5.7 ± 0.08
GAN-extra 53.0 ± 1.18 481.5 ± 5.74 798.9 635.5 1.98 ± 0.02 5.39 ± 0.06 8.8 ± 0.08
CorrDiff-extra 55.2 ± 0.96 438.6 ± 4.85 833.7 787.5 1.91 ± 0.02 4.91 ± 0.05 4.0 ± 0.06
EnScale-extra 36.2 ± 0.8 456.4 ± 4.98 844.8 776.5 1.92 ± 0.02 5.05 ± 0.05 5.5 ± 0.07
EnScale-t-extra 36.2 ± 0.81 453.7 ± 5.08 850.1 790.5 1.94 ± 0.02 5.09 ± 0.05 5.3 ± 0.06
sfcWind
method MSE $\downarrow$ ES $\downarrow$ $_{\text{pred}} ES$$_{\text{var}}$ CRPS $\downarrow$ CRPS-mp $\downarrow$ RALSD $\downarrow$
NN-det-extra 2.6 ± 0.03 196.4 ± 1.01 196.5 0.0 1.13 ± 0.01 1.44 ± 0.01 2.0 ± 0.02
EasyUQ-extra 4.2 ± 0.03 141.2 ± 0.79 255.6 229.0 0.88 ± 0.0 2.22 ± 0.01 8.9 ± 0.03
Analogues-extra 3.9 ± 0.04 180.8 ± 1.49 304.1 246.4 1.13 ± 0.01 1.39 ± 0.01 2.0 ± 0.02
GAN-extra 4.7 ± 0.04 151.5 ± 0.82 269.9 236.9 0.95 ± 0.0 1.16 ± 0.01 3.6 ± 0.03
CorrDiff-extra 3.8 ± 0.04 137.3 ± 0.84 269.5 264.2 0.82 ± 0.0 1.00 ± 0.01 1.4 ± 0.01
EnScale-extra 2.6 ± 0.03 133.1 ± 0.87 264.4 262.7 0.85 ± 0.0 1.05 ± 0.01 1.8 ± 0.02
EnScale-t-extra 2.7 ± 0.03 134.8 ± 0.91 265.1 260.5 0.86 ± 0.0 1.06 ± 0.01 1.8 ± 0.02
rsds
method MSE $\downarrow$ ES $\downarrow$ $_{\text{pred}} ES$$_{\text{var}}$ CRPS $\downarrow$ CRPS-mp $\downarrow$ RALSD $\downarrow$
NN-det-extra 1914.9 ± 33.34 4888.1 ± 40.52 4883.9 0.0 29.89 ± 0.28 25.76 ± 0.22 3.5 ± 0.03
EasyUQ-extra 3217.8 ± 40.21 3520.5 ± 29.43 6613.2 6188.4 23.60 ± 0.2 46.98 ± 0.32 10.3 ± 0.02
Analogues-extra 2560.2 ± 36.13 4277.1 ± 44.87 7221.1 5901.3 28.19 ± 0.22 21.02 ± 0.18 1.8 ± 0.02
GAN-extra 2906.5 ± 40.6 3566.5 ± 31.14 6197.6 5265.8 23.43 ± 0.19 17.70 ± 0.15 3.7 ± 0.03
CorrDiff-extra 3466.7 ± 54.06 3100.6 ± 28.9 6243.2 6285.3 21.85 ± 0.19 15.58 ± 0.15 1.5 ± 0.01
EnScale-extra 1936.0 ± 30.34 3371.1 ± 31.24 6930.1 7118.1 23.29 ± 0.19 21.45 ± 0.16 2.9 ± 0.02
EnScale-t-extra 1989.1 ± 29.1 3392.3 ± 30.77 6995.6 7211.4 23.41 ± 0.2 21.89 ± 0.16 2.9 ± 0.02
A.5 Further time series evaluation
Table A.2: Lag-1 autocorrelations (ACF) for each variablea
Method tas pr sfcWind rsds
NN-det 0.01 0.17 0.08 0.11
EasyUQ 0.03 0.08 0.15 0.03
Analogues 0.1 0.12 0.28 0.04
GAN 0.06 0.1 0.21 0.02
CorrDiff 0.01 0.23 0.1 0.09
EnScale 0.08 0.09 0.22 0.06
EnScale-t 0.01 0.04 0.05 0.02
aWe calculate ACFs in each location for both the samples and the RCM data. We average the absolute error (magnitude of the difference of sampled minus target, regardless of the sign) across locations.
Figure A.14: Power spectral densities
Table A.3: RALSD-avg
method tas pr sfcWind rsds
NN-det 0.466 8.598 1.392 4.050
EasyUQ 5.958 7.610 8.501 9.966
Analogues 0.150 0.344 0.153 0.170
GAN 1.999 6.890 3.183 4.066
CorrDiff 0.394 0.951 0.432 0.752
EnScale 0.579 1.456 0.707 1.293
EnScale-t 0.589 1.429 0.654 1.274
Appendix B Details on methods
B.1 Conditional independence assumptions for multi-step approach
We explain downscaling via coarse correction in Sec. 2.2. Here, we approximate the full distribution by assuming that $Y$ is independent of $X$ given $Z$. Formally, this means $Y\mathrel{\perp\!\!\!\perp}X|Z$. One can show that the concatenation of the coarse model predicting $p_{Z|X}$ with the model for $p_{Y|Z}$ yields $p_{Y|X}$ under this assumption (see proof below). To verify validity of this approximation, we compare the loss when predicting $Y$ from $Z$ only versus from $Z,X$. If $p_{Y|X,Z}\neq p_{Y|Z},$ the loss should decrease in the second case. For simplicity, we checked these losses in a univariate setup, predicting only one variable at a time, and using one dense model as the super-resolution model (without $\tilde{Z}$ as an intermediate step). We checked precipitation and temperature. For precipitation, we could not observe any difference in loss. For temperature, the loss improves by about 2% when adding $X$ as an additional predictor. This indicates that the conditional independence is a close approximation.
The final setup is more complicated. In Sec. 2.4, we additionally describe the splitting of the super-resolution model into several intermediate steps. This requires additional approximations, which we have not verified explicitly.
Under the assumption $Y\mathrel{\perp\!\!\!\perp}X|Z$, the concatenation of coarse model and super-resolution model yields $p_{Y|X}$. We prove it in the following. We use that $A\mathrel{\perp\!\!\!\perp}B|C$ if $p_{A|B,C}=p_{A|C}$ to simplify the full conditional distribution. Here, we assume all distributions to have a density and denote those by small $p$. To model the conditional density $p_{Y|X}$, we integrate the density of the full distribution $p_{Y,Z}$ over all possible $Z$. In the following, we drop the subscripts of the densities for simplicity and use the corresponding lower case letters as arguments, i.e. $p_{Y|X}(y|x)$ will be written as $p(y|x)$.
$p(y|x)$ $=\int p(y,z|x)dz$
$=\int p(y|x,z)p(z|x)dz$
$\mathrel{{\mathop{=}\limits^{(1)}}}\int p(y|z)p(z|x)dz$
where (1) holds because $Y\mathrel{\perp\!\!\!\perp}X|Z$. The integrand in final line is exactly the joint distribution $(Z,Y)|X$ that the concatenation of the two models yields. Then, integrating out $Z$ corresponds to selecting the marginal $Y|X$, equivalent to dropping the components $Z$. Thus, our three-step approach learns $p_{Y|X}$ under the given conditional independence assumptions.
B.2 Conditional independence assumptions for time series generation
Formally, our approach in Sec. 2.3 relies on suitable conditional independence assumptions, similar to the above discussion about the downscaling via coarse correction (App. B.1 ). For simplicity, we omit the intermediate step via $\tilde{Z}$ in this description, and just consider the two steps $X_{t}\rightarrow Z_{t}$ and $Z_{t}\rightarrow Y_{t}$. Here, the key approximation is that we assume $Y_{t}$ to not depend on $Y_{t-1}$ when given $Z_{t}$. In other words, only temporal correlations on the coarse scale (in $Z_{t}$) are considered and further dependencies on the local scale (in $Y_{t}$) are neglected.
We test this approximation by comparing the losses in different prediction setups. We predict $Y_{t}$ from $Z_{t}$ only and from $Z_{t},Y_{t-1}$, checking if the assumption $Y_{t}\mathrel{\perp\!\!\!\perp}Y_{t-1}|Z_{t}$ is valid. For simplicity, we predict one variable at a time, using a single dense model as the super-resolution model. In this case, we test precipitation and surface wind. For precipitation, we find no change in loss between the two setups. For surface wind, including $Y_{t-1}$ improves the loss by about 2%. Like above, this suggests that the conditional independence might not be satisfied perfectly, but is a valid approximation that simplifies computations.
More formally, we require $Y_{t}$ to not depend on $Z_{t-1},Y_{t-1},X_{t}$ when given $Z_{t}$. Similarly, $Z_{t}$ should not depend on $Y_{t-1}$ when given $X_{t},Z_{t-1}$. Thus, the assumptions are:
  1. 1.
    $Y_{t}\mathrel{\perp\!\!\!\perp}(X_{t},Z_{t-1},Y_{t-1})|Z_{t}$
  2. 2.
    $Z_{t}\mathrel{\perp\!\!\!\perp}Y_{t-1}|X_{t},Z_{t-1}$
With these, a similar proof to the above (App. B.1) shows that $p_{Y_{t}|X_{t},Y_{t-1}}$ is indeed learned by our model. As before, we drop subscripts and use the corresponding lower-case letters as arguments.
$p(y_{t}|x_{t},y_{t-1})$ $=\int p(y_{t}|x_{t},y_{t-1},z_{t-1},z_{t})p(z_{t-1},z_{t}|x_{t},y_{t-1})d(z_{t-1},z_{t})$
$=\int p(y_{t}|x_{t},y_{t-1},z_{t-1},z_{t})p(z_{t}|x_{t},y_{t-1},z_{t-1})p(z_{t-1}|x_{t},y_{t-1})d(z_{t-1},z_{t})$
$\mathrel{{\mathop{=}\limits^{(1)}}}\int p(y_{t}|z_{t})p(z_{t}|x_{t},y_{t-1},z_{t-1})p(z_{t-1}|x_{t},y_{t-1})d(z_{t-1},z_{t})$
$\mathrel{{\mathop{=}\limits^{(2)}}}\int p(y_{t}|z_{t})p(z_{t}|x_{t},z_{t-1})p(z_{t-1}|x_{t},y_{t-1})d(z_{t-1},z_{t})$
$=\int p(y_{t}|z_{t})p(z_{t}|x_{t},z_{t-1}(y_{t-1}))dz_{t}$
where in the last step we used that, given $Y_{t-1}$, $Z_{t-1}$ is known, and $Z_{t-1}(Y_{t-1})$ denotes the average pooled $Y_{t-1}$. This last row is what the model generates, with one slight simplification: We predict $\hat{Z}_{t}$ from $\hat{Z}_{t-1}$, instead of the average pooled $\hat{Y}_{t-1}$, such that the time-series generation can happen independently of the super-resolution model.
B.3 Batch-wise approximation
The theoretical expected values (e.g. Eq. (2)) are impossible to calculate. However, during training, they can be replaced by averages over mini-batches of data, yielding an unbiased estimate of the Let $(x_{i},y_{i})\,i\in\{1,...,n\}$ be a mini-batch of data, and $\hat{y_{i}},\hat{y_{i}}^{\prime}\sim Q_{Y|X}(.|x_{i})$ two samples from the conditional distribution of the RCM data given the GCM input $x_{i}$. Then the finite-sample approximation of the loss is
$\hat{\text{Loss}}_{\text{cond}}=\frac{1}{n}\sum_{i=1}^{n}(\lVert y_{i}-\hat{y_{i}}\rVert-\frac{1}{2}\lVert\hat{y_{i}}-\hat{y_{i}}^{\prime}\rVert).$
Note that, per input $x_{i}$, we only use two samples from the generative model, as we found that increasing the batch size has a similar effect as increasing the number of samples. This loss can be minimised via stochastic gradient descent, as usual in machine learning.
B.4 Model architecture
B.4.1 Coarse model
The first coarse model is a dense multilayer perceptron (MLP) and its architecture is as in [Shen2024Engression:Regression]. To enable the model to capture difference between GCM-RCM pairs, we integrate both the index for the GCM and RCM considered: We one-hot-encode GCM and RCM index separately and append both to the input. We kept the GCM-index because our work does not consider extrapolation to unseen GCMs, but preliminary tests suggest that the model performance is very similar also without the GCM-encoding. In addition, the MLP is modified to include random noise that allows for stochastic output. As explained above, the generative model needs to take random initialization as input. Instead of just adding this noise to the input at the beginning, we add the randomness at each layer. We concatenate the hidden units with standard Gaussian noise, where the noise is scaled by a learnable parameter. This concatenation is passed into the next layer. Adding noise in multiple hidden layers allows to learn more complex representations of the Gaussian noise. For example, we find that this yields spatially correlated variability in the output.
Each layer consists of a linear map plus a ReLU activation function. We include skip connections that allow the input to bypass two layers, facilitating gradient flow and improving training stability.
Both the coarse model as well as its temporally-consistent extension consist of 5 hidden layers with sizes 200, 200, 200, 256, 256. The concatenated Gaussian noise has dimension 10 in each layer. We also experimented with larger noise dimension, but found no improvement in performance. We found better performance when adding a first layer to pre-process the inputs: Each input variable is compressed separately with a linear layer and a ReLU-activation, and only afterwards all variables are concatenated. In this way, the model is less sensitive to differences in units or scales in different input variables.
B.4.2 Super-resolution model
[…]
We choose 9 nearest neighbors of the low-resolution input in the first upsampling step. Afterwards, 5 noise channels of Gaussian i.i.d. noise are concatenated. For the stochastic refinement, 25 nearest neighbors in the high-resolution grid are used and linearly aggregated to a latent space of dimension 12 for each pixel. The shared MLP has two hidden layers of hidden dimension 12 each. We add a small weight decay penalty (scaling 1e-3), as we find that this improves image quality.
As for the coarse model, we enable learning RCM-specific details by integrating the index for the GCM-RCM pair into the sparse local layers. In the first linear upsampling step, the learnable weights are pair-specific, i.e. the model can learn a different set of weights for each pair. The stochastic refinement does not depend on the GCM-RCM pair and, thus, learned stochasticity is pair-agnostic. We also tested pair-specific stochastic modeling, but found no improvement in the loss.
B.5 Training
We train all models using the Adam optimizer with learning rate 0.0001. We check convergence of the test loss. The coarse and temporal coarse model are trained for 200 epochs. The sparse layers in the super-resolution model are also trained for the 200 epochs (stages 1–3) and for 100 epochs (last stage).
B.6 Details on benchmarks
B.6.1 Deterministic neural network
We use a similar model architecture to the dense models in our generative approach (see Sec. B.4), using a fully-connected neural network with alternating linear maps and ReLU activation. However, unlike for the generative models, there is no noise input needed for the deterministic benchmark. The layers have sizes 3612, 500, 500, 500, 1024, 4096, 16384.
The data transformations are the same as for EnScale: The predictors are scaled with a simple standardization, and the targets are transformed to uniform. The deterministic network converges very quickly. We use the model checkpoints after 25 epochs. Afterwards, the test loss increases again.
B.6.2 EasyUQ
EasyUQ’s predictions are optimal under a monotonicity assumption, in the sense that they minimize the CRPS. The monotonicity assumption describes that the targets $Y$ tend to increase as the regression model’s predictions $\hat{\mu}$ also increase. This is a natural requirement for a useful statistical model. However, if the regression model is deficient, this assumption might be violated. EasyUQ will still yield predictive distributions, but optimality is not guaranteed in this case. For more details on EasyUQ and its mathematical guarantees, see [Walz2024EasyOutput] and [Henzi2021IsotonicRegression].
We implement EasyUQ based on the Python package isodisreg [WalzIsodisreg]. We train EasyUQ on the training data that was also used to train the NN-det, separately for each location. After training, we retrieve the predicted distributions on the test data and sample from them, independently in each location, time point and for each variable. For this, we slightly adjust the default prediction method from the Python package in order to be able to take in a different set of probabilities for each point in time. These samples can then be compared to the samples from EnScale.
B.7 Analogues
We provide more details on our implementation of the analogues in the following: before generating samples, we calculate a distance matrix capturing pairwise similarities between days in a GCM dataset, including train and test periods. To reduce computational complexity, we restrict comparisons to days within a 50-calendar-day window. For each pair of days, we compute the Euclidean distance of the GCM’s spatial fields separately per variable, considering all four target variables as well as sea level pressure. To aggregate across variables, we follow a rank-based approach: for each variable and day, we sort the distances to the other days, convert them to ranks. Finally, we average the ranks over all variables. This approach standardizes the distances, avoiding scale issues from different units and dependence on data transformations. In summary, two days are considered to be “close” if they are similar w.r.t. all the five GCM variables.
Using this distance matrix, we can approximate $p^{\iota}_{Y|X}$ independently for each GCM-RCM pair. For each day in the test data, we identify the $k=5$ days in the train data with the smallest distance, so-called nearest neighbors (again within a 50-day calendar window). We approximate the conditional $p^{\iota}_{Y|X}$ by the discrete distribution with these $k$ data points. To generate an ensemble, we sample with replacement from these $k$ nearest neighbors.
The choice of $k$ balances the prediction error with the variability: smaller values of $k$ imply closer neighbors and smaller predictions errors but low variability. Larger $k$ increase diversity at the cost of larger prediction error. We found $k=5$ to be a good compromise.
We are aware that a single GCM dataset with about 30k train data points is rather small for an analogue-based approach. However, combining data from multiple GCMs or RCMs gives further issues due to distribution shifts, which is why we include the simple version as a conceptually easy benchmark.
B.8 Proof of variance decomposition
We prove the decomposition from the main text using law of total variance $\operatorname{Var}(Y)=E[\operatorname{Var}(Y|Z)]+\operatorname{Var}(E[Y|Z])$. However, we apply it to $\operatorname{Var}^{\iota}(Y|X)$, s.t. we don’t only condition on $Z$ but also on $X$:
$\operatorname{Var}^{\iota}(Y|X)=E[\operatorname{Var}^{\iota}(Y|Z,X)|X]+\operatorname{Var}^{\iota}(E^{\iota}[Y|Z,X]|X),$
We explained above (see App. B.1) that our approach to downscaling via coarse correction assumes $Y\mathrel{\perp\!\!\!\perp}X|Z$. Hence, $\operatorname{Var}^{\iota}(Y|Z,X)=\operatorname{Var}^{\iota}(Y|Z)$ and $E^{\iota}[Y|Z,X]=E^{\iota}[Y|Z]$, which concludes the proof.